### v1

# Equilibrium-distribution-function based mesoscopic finite-difference methods for partial differential equations: Modeling and Analysis

Abstract
In this paper, based on the idea of direct modeling with equilibrium distribution functions (EDFs) we develop a general framework of mesoscopic numerical method (MesoNM) for macroscopic partial differential equations (PDEs), including but not limited to nonlinear convection-diffusion equation (NCDE) and the Navier-Stokes equations (NSEs). Different from the mesoscopic lattice Boltzmann method, this kind of MesoNM is an EDF-based mesoscopic finite-difference (MesoFD) method, and through taking the moments of MesoFD scheme its macroscopic version, called MMFD method, can be derived directly. Both MesoFD scheme and MMFD schemes are muti-level FD methods, MesoFD scheme being mesoscopic, and MMFD scheme being its macroscopic form which has the form of the central FD scheme. They are unified FD schemes for PDEs and can be implicit or explicit forms as needed. The macroscopic moment equations (MEs) can be derived from the MesoFD or MMFD scheme through the Taylor expansion method, and the common PDEs can be recovered from the MEs by using direct Taylor expansion method. Moreover, the stability of MMFD scheme is analyzed for linear CDE with anisotropic diffusion, and the stability conditions of a two-level explicit MMFD scheme and a two-level $\theta$-MMFD scheme (hybrid explicit and implicit MMFD scheme) are obtained, respectively. Finally, we note that some existing lattice Boltzmann (LB) based macroscopic FD models for the NSEs and NCDE are the special cases of present MMFD, which can be considered as a unified framework of FD schemes for PDEs, from this point of view.
I Introduction
The lattice Boltzmann (LB) method, as an effective mesoscopic numerical method (MesoNM) based on kinetic theory, has gained great success in the study of complex flows, heat and mass transfer described by the Navier-Stokes equations (NSEs), nonlinear convection-diffusion equations (NCDEs), and the coupled systems of them [1, 2, 3, 4]. Up to now, although there are many different types of LB models [16], the most popular ones are the single-relaxation-time LB (SRT-LB) model or lattice Bhatnagar-Gross-Krook (LBGK) model [5, 6], the two-relaxation-time LB (TRT-LB) model [7, 8, 9, 10, 11], and the general or multple-relaxation-time LB (MRT-LB) model [12, 13, 14, 15]. The MRT-LB model has multiple relaxation parameters, and compared to the SRT version, it can achieve more stable and more accurate results through adjusting the free relaxation parameters [12, 13, 14, 15], also the TRT-LB model with only one free relaxation parameter [7, 8, 9, 10, 11]. Actually, the classic MRT-LB model can be more general through introducing a block-lower-triangular-relaxation matrix [16], and rectangular grid [17].
It is known that the LB method is based on the evolution of the distribution function (DF) in all lattice velocity directions at each lattice node, and the macroscopic conserved quantities are computed from the moments of the DF. This evolution based on the DF brings us several challenging problems that need to be addressed: (1) a large amount of memory is required; (2) the initial and boundary conditions for macroscopic quantities should be treated carefully; (3) more importantly, the theoretical analysis (e.g., stability analysis) on the LB method is complicated and difficult. To overcome these problems, it is desirable to explore the connection between LB method and the traditional numerical method. In the past years, some efforts have been made to derive the macroscopic numerical scheme from the LB method.
In the LB community, it is well known that the SRT-LB model with relaxation time $\tau=1$ can be written as a macroscopic FD scheme [18, 19, 20, 21]. In order to overcome the limitation $\tau=1$, the lattice kinetic scheme was proposed in Ref. [20], and several modified versions have been developed in the literature. In addition to those derived from the SRT-LB model [18, 19, 20, 21], up to now there have been many research works on macroscopic FD schemes derived or developed from LB models, e.g., the TRT-LB model [9, 22, 10], and some other type of macroscopic FD schemes, such as the so-called simplified lattice Boltzmann method [23, 24] and its modified models [25, 26], and the recursive FD LB models [27, 28]. In addition, for the one-dimensional case, some three- and four-level FD schemes can also be obtained from the LB models [29, 30, 31]. It should be noted that based on the recurrence equations with $s^{+}+s^{-}=2$, a class of three-level FD schemes can be derived from TRT-LB model [9, 22, 10], which contains the FD scheme obtained by the SRT-LB model as its special case (setting $s^{+}=s^{-}=1$), where $s^{+}$ and $s^{-}$ are the relaxation parameters corresponding to the even-order and odd-order moments, respectively.
It can be seen that the above works on the macroscopic FD schemes [32] derived from the LB models are usually limited to special cases. Recently, Bellotti et al. [33, 34] demonstrated that any LB scheme can be rewritten as a corresponding multi-level FD scheme on the conserved variables, and conducted a rigorous consistency study and the derivation of the modified equations. However, as stated in Ref. [33], the multi-level FD scheme usually has the theoretical value due to its complicated multi-level structure and the formalism based on operators, commutative algebra and polynomials (’we stress once more the fact that recasting a lattice Boltzmann scheme as a Finite Difference scheme on the conserved moments is merely useful from the theoretical perspective, being of moderate interest for actual efficient implementations’ [33]). Based on the results in Refs. [33, 34], some specific LB models and corresponding multi-level FD schemes with high-order accuracy for one- or two-dimensional simple problems have been designed [31, 35]. According to the previous work [16], Chen et al. [36] presented a general propagation multiple-relaxation-time lattice Boltzmann (GPMRT-LB) model, and then following the works [33, 34], they obtained the macroscopic FDS and modified equations corresponding to the GPMRT-LB model.
Different from the above methods for deriving the FD schems from the existing LB models, in this work, we uses the equilibrium distribution function (EDF) to represent the distribution function (DF), and propose a novel class of mesoscopic numerical methods (MesoNMs), called the mesoscopic finite-difference (MesoFD) schemes, for partial differential equations (PDEs) with the source terms. Then the MesoFD based macroscopic finite difference (MMFD) schemes can be directly obtained by taking the moments of the MesoFD scheme. Both of them are of multi-level implicit or explicit structure. The MMFD scheme can be viewed as a unified framework of FD schemes for PDEs, and the commonly used FD schemes derived from LB models and traditional FD schemes are its special cases.
The rest of this paper is organized as follows. In Sec. II, the MesoFD method and the corresponding MMFD scheme are presented. The direct Taylor expansion is adopted to obtain the corresponding moment equations and PDEs of the present MesoFD and MMFD methods, and we present some special cases of the MesoFD and MMFD methods in Sec. III, and followed by some numerical tests in Sec. IV. Finally, some conclusions are summarized in Sec. V.
II Mesoscopic finite difference method and macroscopic finite difference scheme
The MesoFD method comes from the idea of representing the distribution function (DF) or nonequilibrium distribution function (NEDF) with equilibrium distribution function (EDF), which is an extension of the previous work [28].
[…]
$f_{j}(\mathbf{x},t)=\sum_{k,q}\big{[}a_{kq}f_{j}^{eq[k,q]}+\Delta tb_{kq}F_{j}^{[k,q]}\big{]},$ (1)
with
$f_{j}^{eq[k,q]}=f_{j}^{eq}(\mathbf{x}+k\mathbf{c}_{j}\Delta t,t+q\Delta t),\ F_{j}^{[k,q]}=F_{j}(\mathbf{x}+k\mathbf{c}_{j}\Delta t,t+q\Delta t),\ j=0,1,\cdots,b-1,$ (2)
[…]
From Eq. (1) we can also obtain the evolution equations of required moments. Similar to the LB method, the unknown macroscopic conserved variable(s), that is, the zeroth moment $\phi(\mathbf{x},t)$ for scalar PDE, such as NCDE, or the zeroth moment $\rho(\mathbf{x},t)$ and the first moment $(\rho\mathbf{u})(\mathbf{x},t)$ for NSEs, can be computed by
$\phi(\mathbf{x},t)=\sum_{j}f_{j}(\mathbf{x},t),$ (3a)
$\rho(\mathbf{x},t)=\sum_{j}f_{j}(\mathbf{x},t),\quad(\rho\mathbf{u})(\mathbf{x},t)=\sum_{j}\mathbf{c}_{j}f_{j}(\mathbf{x},t).$ (3b)
In this case, the corresponding MMFD scheme can be obtained from Eq. (1),
[…]
$\ \ \ \ \ \ \ \ \ (\rho\mathbf{u})(\mathbf{x},t)=\sum_{j}\sum_{k,q}\mathbf{c}_{j}\big{[}a_{kq}f_{j}^{eq[k,q]}+\Delta tb_{kq}F_{j}^{[k,q]}\big{]}.$ (5b)
Furthermore, through using a divide-and-conquer strategy, we can extend the single-distribution-function scheme Eq. (5) to a double-distribution-function one:
$\rho(\mathbf{x},t)=\sum_{j}\sum_{k,q}\big{[}a_{kq}f_{j}^{eq[k,q]}+\Delta tb_{kq}F_{j}^{[k,q]}\big{]},$ (6a)
$(\rho\mathbf{u})(\mathbf{x},t)=\sum_{j}\sum_{k,q}\mathbf{c}_{j}\big{[}\bar{a}_{kq}f_{j}^{eq[k,q]}+\Delta t\bar{b}_{kq}F_{j}^{[k,q]}\big{]},$ (6b)
in which two different sets of parameters, $\{a_{kq},b_{kq}\}$ and $\{\bar{a}_{kq},\bar{b}_{kq}\}$ are used.
Remark 1 We can replace $f_{j}$ and $a_{00}$ in Eq. (1) by $f_{j}^{ne}$ and $a_{00}-1$, respectively, then we obtain the equivalent form of Eq. (1). Therefore, if the summation is only made for the case $k=q$, one can easily find that the scheme in Ref. [28] is just a special case of the MesoFD scheme (1). Unless otherwise stated, the following discussion is about Eq. (1).
II.1 The central difference representation of MesoFD method
Interestingly, both MesoFD and MMFD schemes can be represented by the central difference format. First, we define the following central difference operators through introducing a function $g(\mathbf{x})$, and two integers $j$ and $k$,
$\Delta_{j,k}^{(1)}g(\mathbf{x})=g(\mathbf{x}+k\mathbf{c}_{j}\Delta t)-g(\mathbf{x}-k\mathbf{c}_{j}\Delta t),$
$\Delta_{j,k}^{(2)}g(\mathbf{x})=g(\mathbf{x}+k\mathbf{c}_{j}\Delta t)-2g(\mathbf{x})+g(\mathbf{x}-k\mathbf{c}_{j}\Delta t),$
$\Delta_{j}^{(1)}g(\mathbf{x})=\Delta_{j,1}^{(1)}g(\mathbf{x})=g(\mathbf{x}+\mathbf{c}_{j}\Delta t)-g(\mathbf{x}-\mathbf{c}_{j}\Delta t),$
$\Delta_{j}^{(2)}g(\mathbf{x})=\Delta_{j,1}^{(2)}g(\mathbf{x})=g(\mathbf{x}+\mathbf{c}_{j}\Delta t)-2g(\mathbf{x})+g(\mathbf{x}-\mathbf{c}_{j}\Delta t).$ (7)
It is easy to obtain
$\Delta_{j,k}^{(1)}=-\Delta_{j,-k}^{(1)}=-\Delta_{\bar{j},k}^{(1)},\quad\Delta_{j,k}^{(2)}=\Delta_{j,-k}^{(2)}=\Delta_{\bar{j},k}^{(2)},$
$\Delta_{0,k}^{(1)}=\Delta_{0,k}^{(2)}=\Delta_{j,0}^{(1)}=\Delta_{j,0}^{(2)}=0,$ (8)
where $\bar{j}$ is the opposite direction of $j$, $\mathbf{c}_{\bar{j}}=-\mathbf{c}_{j}$, $\mathbf{c}_{0}=\mathbf{0}$.
Using the above operators, one can easily obtain the central difference representation of MesoFD scheme (1):
$f_{j}(\mathbf{x},t)=\sum_{k,q}a_{kq}\big{[}1+\frac{1}{2}(\Delta_{j,k}^{(2)}+\Delta_{j,k}^{(1)})\big{]}f_{j}^{eq}(\mathbf{x},t+q\Delta t)+\Delta t\sum_{k,q}b_{kq}\big{[}1+\frac{1}{2}(\Delta_{j,k}^{(2)}+\Delta_{j,k}^{(1)})\big{]}F_{j}(\mathbf{x},t+q\Delta t),$
$f_{\bar{j}}(\mathbf{x},t)=\sum_{k,q}a_{kq}\big{[}1+\frac{1}{2}(\Delta_{j,k}^{(2)}-\Delta_{j,k}^{(1)})\big{]}f_{\bar{j}}^{eq}(\mathbf{x},t+q\Delta t)+\Delta t\sum_{k,q}b_{kq}\big{[}1+\frac{1}{2}(\Delta_{j,k}^{(2)}-\Delta_{j,k}^{(1)})\big{]}F_{\bar{j}}(\mathbf{x},t+q\Delta t).$
For a given distribution function $g_{j}$, let
$g_{j}^{+}=\frac{1}{2}(g_{j}+g_{\bar{j}}),\ g_{j}^{-}=\frac{1}{2}(g_{j}-g_{\bar{j}}),$ (11)
then we have
$g_{j}^{+}=g_{\bar{j}}^{+},\ g_{j}^{-}=-g_{\bar{j}}^{-},\ g_{j}=g_{j}^{+}+g_{j}^{-},\ g_{\bar{j}}=g_{j}^{+}-g_{j}^{-}.$ (12)
With the help of Eqs. (II.1), (11) and (12), the summation and the difference of Eqs. (9) and (10) become
$f_{j}^{+}(\mathbf{x},t)$ $=$ $\sum_{k,q}a_{kq}\big{[}f_{j}^{eq,+}+\frac{1}{2}(\Delta_{j,k}^{(2)}f_{j}^{eq,+}+\Delta_{j,k}^{(1)}f_{j}^{eq,-})\big{]}(\mathbf{x},t+q\Delta t)+\Delta t\sum_{k,q}b_{kq}\big{[}F_{j}^{+}+\frac{1}{2}(\Delta_{j,k}^{(2)}F_{j}^{+}+\Delta_{j,k}^{(1)}F_{j}^{-})\big{]}(\mathbf{x},t+q\Delta t)$ (13a)
$=$ $\sum_{k,q}\big{[}a_{kq}f_{j}^{eq,+}+\Delta tb_{kq}F_{j}^{+}\big{]}(\mathbf{x},t+q\Delta t)+\frac{1}{2}\sum_{q}\sum_{k=1}^{m_{q}}\big{[}a_{kq}^{+}\Delta_{j,k}^{(2)}f_{j}^{eq,+}+a_{kq}^{-}\Delta_{j,k}^{(1)}f_{j}^{eq,-})\big{]}(\mathbf{x},t+q\Delta t)$
$+\frac{\Delta t}{2}\sum_{q}\sum_{k=1}^{m_{q}}\big{[}b_{kq}^{+}\Delta_{j,k}^{(2)}F_{j}^{+}+b_{kq}^{-}\Delta_{j,k}^{(1)}F_{j}^{,-})\big{]}(\mathbf{x},t+q\Delta t),$
$f_{j}^{-}(\mathbf{x},t)$ $=$ $\sum_{k,q}a_{kq}\big{[}f_{j}^{eq,-}+\frac{1}{2}(\Delta_{j,k}^{(2)}f_{j}^{eq,-}+\Delta_{j,k}^{(1)}f_{j}^{eq,+})\big{]}(\mathbf{x},t+q\Delta t)+\Delta t\sum_{k,q}b_{kq}\big{[}F_{j}^{-}+\frac{1}{2}(\Delta_{j,k}^{(2)}F_{j}^{-}+\Delta_{j,k}^{(1)}F_{j}^{+})\big{]}(\mathbf{x},t+q\Delta t),$ (13b)
$=$ $\sum_{k,q}\big{[}a_{kq}f_{j}^{eq,-}+\Delta tb_{kq}F_{j}^{-}\big{]}(\mathbf{x},t+q\Delta t)+\frac{1}{2}\sum_{q}\sum_{k=1}^{m_{q}}\big{[}a_{kq}^{+}\Delta_{j,k}^{(2)}f_{j}^{eq,-}+a_{kq}^{-}\Delta_{j,k}^{(1)}f_{j}^{eq,+})\big{]}(\mathbf{x},t+q\Delta t)$
$+\frac{\Delta t}{2}\sum_{q}\sum_{k=1}^{m_{q}}\big{[}b_{kq}^{+}\Delta_{j,k}^{(2)}F_{j}^{-}+b_{kq}^{-}\Delta_{j,k}^{(1)}F_{j}^{,+})\big{]}(\mathbf{x},t+q\Delta t),$
where
$a_{kq}^{+}=a_{kq}+a_{-kq},~{}a_{kq}^{-}=a_{kq}-a_{-kq},~{}b_{kq}^{+}=b_{kq}+b_{-kq},~{}b_{kq}^{-}=b_{kq}-b_{-kq}.$ (14)
Based on the following equations,
$\sum_{j}h_{j}^{+}=\sum_{j}h_{j},\sum_{j}h_{j}^{-}=0,~{}\sum_{j}\mathbf{c}_{j}h_{j}^{+}=\mathbf{0},~{}\sum_{j}\mathbf{c}_{j}h_{j}^{-}=\sum_{j}\mathbf{c}_{j}h_{j},~{}h_{j}\in\{f_{j},f_{j}^{eq},F_{j}\},$ (15)
we can take the zeroth moment of $f_{j}^{+}$ and the first moment of $f_{j}^{-}$ in Eq. (13), respectively, and get the MMFD schemes for the NCDE and NSEs, which are equivalent to those in Eqs. (4) and (6)
$\phi(\mathbf{x},t)$ $=$ $\sum_{k,q}\big{[}a_{kq}\phi+\Delta tb_{kq}S\big{]}(\mathbf{x},t+q\Delta t)+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}\big{[}a_{kq}^{+}\Delta_{j,k}^{(2)}f_{j}^{eq,+}+\Delta tb_{kq}^{+}\Delta_{j,k}^{(2)}F_{j}^{+}\big{]}(\mathbf{x},t+q\Delta t)$ (16)
$+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}\big{[}a_{kq}^{-}\Delta_{j,k}^{(1)}f_{j}^{eq,-}+\Delta tb_{kq}^{-}\Delta_{j,k}^{(1)}F_{j}^{-}\big{]}(\mathbf{x},t+q\Delta t);$
$\rho(\mathbf{x},t)$ $=$ $\sum_{k,q}\big{[}a_{kq}\rho+\Delta tb_{kq}S\big{]}(\mathbf{x},t+q\Delta t)+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}\big{[}a_{kq}^{+}\Delta_{j,k}^{(2)}f_{j}^{eq,+}+\Delta tb_{kq}^{+}\Delta_{j,k}^{(2)}F_{j}^{+}\big{]}(\mathbf{x},t+q\Delta t)$ (17a)
$+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}\big{[}a_{kq}^{-}\Delta_{j,k}^{(1)}f_{j}^{eq,-}+\Delta tb_{kq}^{-}\Delta_{j,k}^{(1)}F_{j}^{-}\big{]}(\mathbf{x},t+q\Delta t),$
$(\rho\mathbf{u})(\mathbf{x},t)$ $=$ $\sum_{k,q}\big{[}\bar{a}_{kq}(\rho\mathbf{u})+\Delta t\bar{b}_{kq}\mathbf{F}\big{]}(\mathbf{x},t+q\Delta t)+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}\mathbf{c}_{j}\big{[}\bar{a}_{kq}^{+}\Delta_{j,k}^{(2)}f_{j}^{eq,-}+\Delta t\bar{b}_{kq}^{-}\Delta_{j,k}^{(2)}F_{j}^{-}\big{]}(\mathbf{x},t+q\Delta t)$ (17b)
$+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}\mathbf{c}_{j}\big{[}\bar{a}_{kq}^{-}\Delta_{j,k}^{(1)}f_{j}^{eq,+}+\Delta t\bar{b}_{kq}^{-}\Delta_{j,k}^{(1)}F_{j}^{+}\big{]}(\mathbf{x},t+q\Delta t),$
where $S,\mathbf{F}$ in Eqs. (16) and (17) and the parameters in Eq. (17b) are defined as $S=\sum_{j}F_{j}=\sum_{j}F_{j}^{+},~{}\mathbf{F}=\sum_{j}\mathbf{c}_{j}F_{j}=\sum_{j}\mathbf{c}_{j}F_{j}^{-},~{}$ $\bar{a}_{kq}^{+}=\bar{a}_{kq}+\bar{a}_{-kq},~{}\bar{a}_{kq}^{-}=\bar{a}_{kq}-\bar{a}_{-kq},~{}\bar{b}_{kq}^{+}=\bar{b}_{kq}+\bar{b}_{-kq},~{}\bar{b}_{kq}^{-}=\bar{b}_{kq}-\bar{b}_{-kq}$, and $\bar{a}_{kq},\bar{b}_{kq}$ are used as those in Eq. (6). If needed, one can take the higher-order moments in Eq. (13) to get the evolution equations of these higher-order moments.
It can be seen from Eqs. (16) and (17) that the MMFD scheme consists of central difference schemes on multiple-time levels. Based on the MesoFD scheme (Eq. (1)) and the MMFD scheme (Eq. (16) or Eq. (17)), we can derive the macroscopic PDEs from mesoscopic and macroscopic ways, respectively.
II.2 The two-level central MMFD scheme
As we know, the two-level FD scheme is the most commonly used one for its efficiency . In this part, we only consider the two-level central MMFD scheme for NCDE, and the two-level central MMFD scheme for NSEs can be given similarly. Taking $q=-1$ in Eq. (16), we obtain the explicit MMFD (eMMFD) scheme
$\phi(\mathbf{x},t)$ $=$ $\sum_{k}\big{[}a_{k,-1}\phi+\Delta tb_{k,-1}S\big{]}(\mathbf{x},t-\Delta t)+\frac{1}{2}\sum_{j}\sum_{k=1}^{m_{-1}}\big{[}a_{k,-1}^{+}\Delta_{j,k}^{(2)}f_{j}^{eq,+}+\Delta tb_{k,-1}^{+}\Delta_{j,k}^{(2)}F_{j}^{+}\big{]}(\mathbf{x},t-\Delta t)$ (18)
$+\frac{1}{2}\sum_{j}\sum_{k=1}^{m_{-1}}\big{[}a_{k,-1}^{-}\Delta_{j,k}^{(1)}f_{j}^{eq,-}+\Delta tb_{k,-1}^{-}\Delta_{j,k}^{(1)}F_{j}^{-}\big{]}(\mathbf{x},t-\Delta t),$
and taking $q\in\{0,-1\}$ in Eq. (16), the implicit MMFD (iMMFD) scheme can be obtained as follows
$\phi(\mathbf{x},t)$ $=$ $\sum_{k}\big{[}a_{k,-1}\phi+\Delta tb_{k,-1}S\big{]}(\mathbf{x},t-\Delta t)+\frac{1}{2}\sum_{j}\sum_{k=1}^{m_{0}}\big{[}a_{k,0}^{+}\Delta_{j,k}^{(2)}f_{j}^{eq,+}+\Delta tb_{k,0}^{+}\Delta_{j,k}^{(2)}F_{j}^{+}\big{]}(\mathbf{x},t)$ (19)
$+\frac{1}{2}\sum_{j}\sum_{k=1}^{m_{0}}\big{[}a_{k,0}^{-}\Delta_{j,k}^{(1)}f_{j}^{eq,-}+\Delta tb_{k,0}^{-}\Delta_{j,k}^{(1)}F_{j}^{-}\big{]}(\mathbf{x},t),$
The eMMFD and iMMFD schemes can be combined to give a more general two-level $\theta$-scheme,
$\phi(\mathbf{x},t)=\sum_{k}\big{[}a_{k}\phi+\Delta tb_{k}S\big{]}(\mathbf{x},t-\Delta t)$
$+\frac{1}{2}\sum_{j}\sum_{k=1}^{m_{0}}\big{[}a_{k}^{+}\Delta_{j,k}^{(2)}(\theta f_{j}^{eq,+}(\mathbf{x},t-\Delta t)+\theta_{1}f_{j}^{eq,+}(\mathbf{x},t))+\Delta tb_{k}^{+}\Delta_{j,k}^{(2)}(\theta F_{j}^{+}(\mathbf{x},t-\Delta t)+\theta_{1}F_{j}^{+}(\mathbf{x},t))\big{]}$
$+\frac{1}{2}\sum_{j}\sum_{k=1}^{m_{0}}\big{[}a_{k}^{-}\Delta_{j,k}^{(1)}(\theta f_{j}^{eq,-}(\mathbf{x},t-\Delta t)+\theta_{1}f_{j}^{eq,-}(\mathbf{x},t))+\Delta tb_{k}^{-}\Delta_{j,k}^{(1)}(\theta F_{j}^{-}(\mathbf{x},t-\Delta t)+\theta_{1}F_{j}^{-}(\mathbf{x},t))\big{]},$ (20)
where $\theta\in[0,1],~{}\theta_{1}=1-\theta,~{}a_{k}^{+}=a_{k}+a_{-k},~{}a_{k}^{-}=a_{k}-a_{-k},$ and $b_{k}^{+}=b_{k}+b_{-k},b_{k}^{-}=b_{k}-b_{-k}$.
III The moment equations of MesoFD method: Direct Taylor expansion
There are four basic analysis methods that can be used to recover the macroscopic NSEs and NCDE from the LB models, i.e., the Chapman-Enskog (CE) analysis [37], the Maxwell iteration (MI) method [38], the direct Taylor expansion (DTE) method [16] and the recurrence equations (RE) method [10, 9, 11]. In Ref. [16], these four methods have been compared, and it is shown that they can give the same equations at the second-order of expansion parameters, while the DTE method is much simpler. The MI method and RE method may not be suitable for analyzing the MesoFD scheme. In what follows, the DTE method is used to analyze the MesoFD scheme at the acoustic scaling.
Applying the Taylor expansion to Eq. (1), one can get
$f_{j}=\sum_{l=0}^{N}\frac{\Delta t^{l}}{l!}\sum_{k,q}\big{[}a_{kq}D_{k,q}^{l}f_{j}^{eq}+\Delta tb_{kq}D_{k,q}^{l}F_{j}\big{]}+O(\Delta t^{N+1}),~{}N\geq 1,$ (21)
where $D_{k,q}=q\partial_{t}+k\nabla\cdot\mathbf{c}_{j}$.
Let
$A_{lm}=\frac{1}{l!}C_{l}^{m}\sum_{k,q}a_{kq}q^{l-m}k^{m},~{}B_{lm}=\frac{1}{l!}C_{l}^{m}\sum_{k,q}b_{kq}q^{l-m}k^{m},$ (22)
where $C_{l}^{m}=\frac{l!}{m!(l-m)!}$. From Eq. (21), we can derive the following equations,
$f_{j}=\sum_{l=0}^{N}\Delta t^{l}\sum_{m=0}^{l}\big{[}A_{lm}\partial_{t}^{l-m}(\nabla\cdot\mathbf{c}_{j})^{m}f_{j}^{eq}+\Delta tB_{lm}\partial_{t}^{l-m}(\nabla\cdot\mathbf{c}_{j})^{m}F_{j}\big{]}+O(\Delta t^{N+1}),~{}N\geq 1,$ (23)
where the following formula is used
$D_{k,q}^{l}=(q\partial_{t}+k\nabla\cdot\mathbf{c}_{j})^{l}=\sum_{m=0}^{l}C_{l}^{m}(q\partial_{t})^{l-m}(k\nabla\cdot\mathbf{c}_{j})^{m}.$ (24)
Taking $N=$1 and 2 in Eq. (23), one can obtain
$f_{j}=A_{00}f_{j}^{eq}+\Delta t(A_{10}\partial_{t}+A_{11}\nabla\cdot\mathbf{c}_{j})f_{j}^{eq}+\Delta tB_{00}F_{j}+O(\Delta t^{2}),$ (25a)
and
$f_{j}$ $=$ $A_{00}f_{j}^{eq}+\Delta t(A_{10}\partial_{t}+A_{11}\nabla\cdot\mathbf{c}_{j})f_{j}^{eq}+\Delta t^{2}\big{[}A_{20}\partial_{t}^{2}+A_{21}\partial_{t}\nabla\cdot\mathbf{c}_{j}+A_{22}(\nabla\cdot\mathbf{c}_{j})^{2}\big{]}f_{j}^{eq}$ (25b)
$\Delta tB_{00}F_{j}+\Delta t^{2}(B_{10}\partial_{t}+B_{11}\nabla\cdot\mathbf{c}_{j})F_{j}+O(\Delta t^{3}),$
respectively.
III.1 The derivation of the moment equation for PDE using DTE method
Based on Eq. (23), we can first derive the moment equation for PDE from the MesoFD method (1). To simplify the following analysis, we denote the moments of $f_{j}$, $f_{j}^{eq}$, and $F_{j}$ as
$M_{0}=\sum_{j}f_{j}^{eq}=\sum_{j}f_{j},\ \ \mathbf{M}_{1}=\sum_{j}\mathbf{c}_{j}f_{j}^{eq},\ \ \mathbf{M}_{k}=\sum_{j}\mathbf{c}_{j}\cdot\cdot\cdot\mathbf{c}_{j}f_{j}^{eq},~{}k>1,$ (26a)
$M_{0F}=\sum_{j}F_{j},\ \ \mathbf{M}_{1F}=\sum_{j}\mathbf{c}_{j}F_{j},\ \ \mathbf{M}_{kF}=\sum_{j}\mathbf{c}_{j}\cdot\cdot\cdot\mathbf{c}_{j}F_{j},~{}k>1,$ (26b)
where $\mathbf{M}_{k}$ and $\mathbf{M}_{kF}$ ($k\geq 0$) are the $k$-th moments of $f_{j}^{eq}$ and $F_{j}$, respectively. Additionally, from Eq. (26a) one can derive the following moment of nonequilibrium distribution function,
$M_{0}^{ne}=\sum_{j}f_{j}^{ne}=\sum_{j}(f_{j}-f_{j}^{eq})=0.$ (27)
Summing Eq. (23) and adopting Eq. (26), one can obtain
$M_{0}$ $=$ $\sum_{l=0}^{N}\Delta t^{l}\sum_{m=0}^{l}\big{[}A_{lm}\partial_{t}^{l-m}(\nabla^{m}\otimes\mathbf{M}_{m})+\Delta tB_{lm}\partial_{t}^{l-m}(\nabla^{m}\otimes\mathbf{M}_{mF})\big{]}+O(\Delta t^{N+1})$
$=$ $A_{00}M_{0}+\sum_{l=1}^{N}\Delta t^{l}\sum_{m=0}^{l}\big{[}A_{lm}\partial_{t}^{l-m}(\nabla^{m}\otimes\mathbf{M}_{m})\big{]}+\sum_{l=0}^{N-1}\Delta t^{l+1}\sum_{m=0}^{l}\big{[}B_{lm}\partial_{t}^{l-m}(\nabla^{m}\otimes\mathbf{M}_{mF})\big{]}+O(\Delta t^{N+1}),~{}N\geq 1,$
where
$\nabla^{m}\otimes\mathbf{M}_{m}=\sum_{j}(\nabla\cdot\mathbf{c}_{j})^{m}f_{j}^{eq}=(\nabla\cdot\cdot\cdot\nabla)\otimes\sum_{j}\mathbf{c}_{j}\cdot\cdot\cdot\mathbf{c}_{j}f_{j}^{eq}.$ (29)
Taking $A_{00}=1$ in Eq. (III.1), we obtain a general moment equation for PDE of order $N$,
$\sum_{l=1}^{N}\Delta t^{l-1}\sum_{m=0}^{l}\big{[}A_{lm}\partial_{t}^{l-m}(\nabla^{m}\otimes\mathbf{M}_{m})\big{]}+\sum_{l=0}^{N-1}\Delta t^{l}\sum_{m=0}^{l}\big{[}B_{lm}\partial_{t}^{l-m}(\nabla^{m}\otimes\mathbf{M}_{mF})\big{]}=O(\Delta t^{N}),~{}N\geq 1,$ (30)
which gives for $N=1,2$, and $3$ that
$A_{10}\partial_{t}M_{0}+A_{11}\nabla\cdot\mathbf{M}_{1}+B_{00}M_{0F}=O(\Delta t),$ (31a)
$A_{10}\partial_{t}M_{0}+A_{11}\nabla\cdot\mathbf{M}_{1}+B_{00}M_{0F}+\Delta t\big{[}A_{20}\partial_{t}^{2}M_{0}+A_{21}\partial_{t}\nabla\cdot\mathbf{M}_{1}+A_{22}\nabla\nabla:\mathbf{M}_{2}\big{]}$
$+\Delta t(B_{10}\partial_{t}M_{0F}+B_{11}\nabla\cdot\mathbf{M}_{1F})=O(\Delta t^{2}),$ (31b)
$A_{10}\partial_{t}M_{0}+A_{11}\nabla\cdot\mathbf{M}_{1}+B_{00}M_{0F}+\Delta t\big{[}A_{20}\partial_{t}^{2}M_{0}+A_{21}\partial_{t}\nabla\cdot\mathbf{M}_{1}+A_{22}\nabla\nabla:\mathbf{M}_{2}\big{]}$
$+\Delta t(B_{10}\partial_{t}M_{0F}+B_{11}\nabla\cdot\mathbf{M}_{1F})+\Delta t^{2}\big{[}A_{30}\partial_{t}^{3}M_{0}+A_{31}\partial_{t}^{2}\nabla\cdot\mathbf{M}_{1}+A_{32}\partial_{t}\nabla\nabla:\mathbf{M}_{2}+A_{33}\nabla^{3}\otimes\mathbf{M}_{3}\big{]}$
$+\Delta t^{2}(B_{20}\partial_{t}^{2}M_{0F}+B_{21}\partial_{t}\nabla\cdot\mathbf{M}_{1F}+B_{22}\nabla\nabla:\mathbf{M}_{2F})=O(\Delta t^{3}),$ (31c)
where
$A_{00}=\sum_{k,q}a_{kq}=1,~{}A_{10}=\sum_{k,q}q\cdot a_{kq},~{}A_{11}=\sum_{k,q}k\cdot a_{kq},$
$A_{20}=\frac{1}{2}\sum_{k,q}q^{2}\cdot a_{kq},~{}A_{21}=\sum_{k,q}q\cdot k\cdot a_{kq},~{}A_{22}=\frac{1}{2}\sum_{k,q}k^{2}\cdot a_{kq},$
$A_{30}=\frac{1}{6}\sum_{k,q}q^{3}\cdot a_{kq},~{}A_{31}=\frac{1}{2}\sum_{k,q}q^{2}\cdot k\cdot a_{kq},~{}A_{32}=\frac{1}{2}\sum_{k,q}q\cdot k^{2}\cdot a_{kq},~{}A_{33}=\frac{1}{6}\sum_{k,q}k^{3}\cdot a_{kq},$
$B_{00}=\sum_{k,q}b_{kq},~{}B_{10}=\sum_{k,q}q\cdot b_{kq},~{}B_{11}=\sum_{k,q}k\cdot b_{kq},$
$B_{20}=\frac{1}{2}\sum_{k,q}q^{2}\cdot b_{kq},~{}B_{21}=\sum_{k,q}q\cdot k\cdot b_{kq},~{}B_{22}=\frac{1}{2}\sum_{k,q}k^{2}\cdot b_{kq}.$ (32)
We note that Eq. (31c) is the moment equation corresponding to a third-order PDE, while Eq. (31b) is a second-order approximation and is also a moment equation for second-order PDE. Eq. (31a) can be taken as their first-order approximation. Some popular second-order PDEs, such as NCDE (DE), convection equation, Poisson equation, and even wave equation can be derived from Eq. (31b) or Eq. (31c) when $A_{lm},B_{lm},\mathbf{\mathbf{M}}_{k}$, and $\mathbf{M}_{kF}$ are properly given.
(1) NCDE
When $M_{0}$ and $\mathbf{M}_{1}$ are not constants, taking $A_{10}=A_{11}$ ($\neq 0$) in Eq. (31), we have
$\partial_{t}M_{0}+\nabla\cdot\mathbf{M}_{1}=-\frac{B_{00}}{A_{10}}M_{0F}+O(\Delta t),$ (33a)
$\partial_{t}M_{0}+\nabla\cdot\mathbf{M}_{1}+\frac{\Delta t}{A_{10}}\big{[}A_{20}\partial_{t}^{2}M_{0}+A_{21}\partial_{t}\nabla\cdot\mathbf{M}_{1}+A_{22}\nabla\nabla:\mathbf{M}_{2}\big{]}$
$=-\frac{B_{00}}{A_{10}}M_{0F}-\frac{\Delta t}{A_{10}}(B_{10}\partial_{t}M_{0F}+B_{11}\nabla\cdot\mathbf{M}_{1F})+O(\Delta t^{2}).$ (33b)
Let
$M_{0F}=S,A_{10}=-B_{00},$ (34)
then we obtain $-\frac{B_{00}}{A_{10}}M_{0F}=S$, and Eq. (33) becomes
$\partial_{t}M_{0}+\nabla\cdot\mathbf{M}_{1}=S+O(\Delta t),$ (35a)
$\partial_{t}M_{0}+\nabla\cdot\mathbf{M}_{1}+\frac{\Delta t}{A_{10}}\big{[}A_{20}\partial_{t}^{2}M_{0}+A_{21}\partial_{t}\nabla\cdot\mathbf{M}_{1}+A_{22}\nabla\nabla:\mathbf{M}_{2}\big{]}$
$=S-\frac{\Delta t}{A_{10}}(B_{10}\partial_{t}S+B_{11}\nabla\cdot\mathbf{M}_{1F})+O(\Delta t^{2}).$ (35b)
From Eq. (35a), we have
$\partial_{t}^{2}M_{0}=-\partial_{t}\nabla\cdot\mathbf{M}_{1}+\partial_{t}S+O(\Delta t),$ (36)
then inserting Eq. (36) into Eq. (35b) yields
$\partial_{t}M_{0}+\nabla\cdot\mathbf{M}_{1}+\frac{\Delta t}{A_{10}}(B_{10}+A_{20})\partial_{t}S$
$=S-\frac{\Delta t}{A_{10}}\big{[}(A_{21}-A_{20})\partial_{t}\nabla\cdot\mathbf{M}_{1}+B_{11}\nabla\cdot\mathbf{M}_{1F}+A_{22}\nabla\nabla:\mathbf{M}_{2}\big{]}+O(\Delta t^{2}).$ (37)
If we set
$(B_{10}+A_{20})\partial_{t}S=0,$ (38)
the moment equation at the leading order of $O(\Delta t^{2})$ can be derived from Eq. (III.1),
$\partial_{t}M_{0}+\nabla\cdot\mathbf{M}_{1}=S-\frac{\Delta t}{A_{10}}\big{[}(A_{21}-A_{20})\partial_{t}\nabla\cdot\mathbf{M}_{1}+B_{11}\nabla\cdot\mathbf{M}_{1F}+A_{22}\nabla\nabla:\mathbf{M}_{2}\big{]}+O(\Delta t^{2}).$ (39)
For a specific NCDE, we can correctly recover the NCDE from Eq. (39) through choosing the moments and parameters. For example, taking
$M_{0}=\phi,~{}\mathbf{M}_{1}=\mathbf{B},~{}\mathbf{M}_{2}=\beta c_{s}^{2}\mathbf{D}+\mathbf{C},~{}\mathbf{M}_{1F}=-\frac{1}{B_{11}}\big{[}(A_{21}-A_{20})\partial_{t}\mathbf{B}+A_{22}\nabla\cdot\mathbf{\mathbf{C}}\big{]},~{}\alpha=-\frac{A_{22}}{A_{10}}\Delta t\beta c_{s}^{2}$ (40)
in Eq. (39) leads to the general NCDE [16, 17] with a constant diffusion coefficient $\alpha$
$\partial_{t}\phi+\nabla\cdot\mathbf{B}=S+\alpha\nabla\nabla:\mathbf{D}+O(\Delta t^{2})=S+\nabla\cdot[\alpha\nabla\cdot\mathbf{D}]+O(\Delta t^{2}).$ (41)
(2) Poisson-type equation or steady CDE [39]
Set $M_{0}=0$ or consider a steady case, Eq. (31b) becomes
$A_{11}\nabla\cdot\mathbf{M}_{1}+\Delta tA_{22}\nabla\nabla:\mathbf{M}_{2}+B_{00}M_{0F}+\Delta tB_{11}\nabla\cdot\mathbf{M}_{1F}=O(\Delta t^{2}).$ (42)
Taking $\mathbf{M}_{1}=\mathbf{B},~{}\mathbf{M}_{2}=\beta c_{s}^{2}\mathbf{D},~{}M_{0F}=S,~{}\mathbf{M}_{1F}=\mathbf{0},~{}B_{00}=-A_{11}$, and $\alpha=-\frac{A_{22}}{A_{11}}\Delta t\beta c_{s}^{2}$ in Eq. (42), the steady CDE is recovered
$\nabla\cdot\mathbf{B}=\alpha\nabla\nabla:\mathbf{D}+S=O(\Delta t^{2}).$ (43)
(3) Wave equation without source term.
Let $A_{20}=-1,~{}A_{lm}=0$ for $(l,m)\neq(2,0)$ and $(2,2)$, and $B_{lm}=0$, for all $l,m$ in Eq. (31c) a wave equation is recovered as follows.
$\partial_{t}^{2}M_{0}=A_{22}\nabla\nabla:\mathbf{M}_{2}+O(\Delta t^{2}).$ (44)
Taking $M_{0}=\phi,~{}\mathbf{M}_{2}=\beta c_{s}^{2}\mathbf{D}$ in Eq. (44), we obtain the following wave equation
$\partial_{t}^{2}\phi=\alpha\nabla\nabla:\mathbf{D}+O(\Delta t^{2})$ (45)
with $\alpha=A_{22}\beta c_{s}^{2}$.
Remark 1: (1) When $\mathbf{M}_{1}=\mathbf{B}=\mathbf{0}$, Eq. (41) becomes a diffusion equation, and we choose $\mathbf{C}=0$, thus $\mathbf{M}_{1F}=0$. In this case, there is no restriction on $A_{11}$, and only $A_{10}\neq 0$ is needed.
(2) The source term $S=0$ or $S$ is a nonzero constant, then Eq. (38) is satisfied. Otherwise, $B_{10}+A_{20}=0$ is needed.
(3) When $\mathbf{D}=\mathbf{0}$, Eq. (41) becomes a convection equation (CE), which means that the CE can also be recovered directly from the MesoFD method. Another way to derive CE from the MesoFD method is to modify the diffusion coefficient $\alpha$ as a small quantity of order $\Delta t$ or $\Delta x$.
(4) If $\mathbf{M}_{1}=0$, and $\mathbf{M}_{3}=0$ in Eq. (31c), then to obtain Eq. (44), $A_{11},A_{21}$ and $A_{33}$ can be non-zero.
(5) If $f_{j}(\mathbf{x},t)$ in Eq. (1) is replaced by $f_{j}(\mathbf{x},t+\Delta t)$, then $A_{l0}=\frac{1}{l!}\sum_{k,q}a_{kq}\cdot q^{l}$ ($l\geq 1$) in Eqs. (22) and (III.1) for the moment equations (30) and (31) should be replaced by $A_{l0}=\frac{1}{l!}\big{[}\sum_{k,q}a_{kq}\cdot q^{l}-1\big{]}$ ($l\geq 1$).
III.2 The MesoFD and MMFD schemes with a three-level explicit structure for PDE
As we know, the LB method for NCDE with a source term is commonly a three-level explicit scheme without correction term, or a two-level implicit scheme with the correction term. We now consider a three-level explicit scheme derived from Eq. (1),
$f_{j}(\mathbf{x},t+\Delta t)=\sum_{q=-1}^{0}\sum_{k=-m_{q}}^{m_{q}}\big{[}a_{kq}f_{j}^{eq[k,q]}+\Delta tb_{kq}F_{j}^{[k,q]}\big{]},$ (46)
and the macroscopic form (MMFDS) can be given as
$\phi(\mathbf{x},t+\Delta t)=\sum_{j}\sum_{q=-1}^{0}\sum_{k=-m_{q}}^{m_{q}}\big{[}a_{kq}f_{j}^{eq[k,q]}+\Delta tb_{kq}F_{j}^{[k,q]}\big{]}.$ (47)
For simplicity, take $k\in\{-1,0,1\}$, and let $a_{k0}=a_{k},a_{k,-1}=\hat{a}_{k},b_{k0}=b_{k},b_{k,-1}=\hat{b}_{k}$, then Eqs. (46) and (47) become
$f_{j}(\mathbf{x},t+\Delta t)$ $=$ $a_{1}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)+a_{0}f_{j}^{eq}(\mathbf{x},t)+a_{-1}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)$ (48)
$+\hat{a}_{1}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t-\Delta t)+\hat{a}_{0}f_{j}^{eq}(\mathbf{x},t-\Delta t)+\hat{a}_{-1}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)$
$+\Delta t\big{[}b_{1}F_{j}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)+b_{0}F_{j}(\mathbf{x},t)+b_{-1}F_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)\big{]}$
$+\Delta t\big{[}\hat{b}_{1}F_{j}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t-\Delta t)+\hat{b}_{0}F_{j}(\mathbf{x},t-\Delta t)+\hat{b}_{-1}F_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)\big{]},$
and
$\phi(\mathbf{x},t+\Delta t)$ $=$ $\sum_{j}\big{[}a_{1}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)+a_{0}f_{j}^{eq}(\mathbf{x},t)+a_{-1}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)\big{]}$ (49)
$+\sum_{j}\big{[}\hat{a}_{1}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t-\Delta t)+\hat{a}_{0}f_{j}^{eq}(\mathbf{x},t-\Delta t)+\hat{a}_{-1}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)\big{]}$
$+\Delta t\sum_{j}\big{[}b_{1}F_{j}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)+b_{0}F_{j}(\mathbf{x},t)+b_{-1}F_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)\big{]}$
$+\Delta t\sum_{j}\big{[}\hat{b}_{1}F_{j}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t-\Delta t)+\hat{b}_{0}F_{j}(\mathbf{x},t-\Delta t)+\hat{b}_{-1}F_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)\big{]}.$
Using Eq. (16), we can rewrite Eq. (49) into another form,
$\phi(\mathbf{x},t+\Delta t)=A_{0}\phi(\mathbf{x},t)+\hat{A}_{0}\phi(\mathbf{x},t-\Delta t)+\Delta t\big{[}B_{0}S(\mathbf{x},t)+\hat{B}_{0}S(\mathbf{x},t-\Delta t)\big{]}$
$+\frac{1}{2}\sum_{j}\big{[}a_{1}^{+}\Delta_{j}^{(2)}f_{j}^{eq,+}+\Delta tb_{1}^{+}\Delta_{j}^{(2)}F_{j}^{+}\big{]}(\mathbf{x},t)+\frac{1}{2}\sum_{j}\big{[}\hat{a}_{1}^{+}\Delta_{j}^{(2)}f_{j}^{eq,+}+\Delta t\hat{b}_{1}^{+}\Delta_{j}^{(2)}F_{j}^{+}\big{]}(\mathbf{x},t-\Delta t)$
$+\frac{1}{2}\sum_{j}\big{[}a_{1}^{-}\Delta_{j}^{(1)}f_{j}^{eq,-}+\Delta tb_{1}^{-}\Delta_{j}^{(1)}F_{j}^{-}\big{]}(\mathbf{x},t)+\frac{1}{2}\sum_{j}\big{[}\hat{a}_{1}^{-}\Delta_{j}^{(1)}f_{j}^{eq,-}+\Delta t\hat{b}_{1}^{-}\Delta_{j}^{(1)}F_{j}^{-}\big{]}(\mathbf{x},t-\Delta t),$ (50)
where
$A_{0}=a_{0}+a_{1}+a_{-1},~{}\hat{A}_{0}=\hat{a}_{0}+\hat{a}_{1}+\hat{a}_{-1},~{}B_{0}=b_{0}+b_{1}+b_{-1},~{}\hat{B}_{0}=\hat{b}_{0}+\hat{b}_{1}+\hat{b}_{-1},$
$a_{1}^{+}=a_{1}+a_{-1},a_{1}^{-}=a_{1}-a_{-1},b_{1}^{+}=b_{1}+b_{-1},b_{1}^{-}=b_{1}-b_{-1},$
$\hat{a}_{1}^{+}=\hat{a}_{1}+\hat{a}_{-1},\hat{a}_{1}^{-}=\hat{a}_{1}-\hat{a}_{-1},\hat{b}_{1}^{+}=\hat{b}_{1}+\hat{b}_{-1},\hat{b}_{1}^{-}=\hat{b}_{1}-\hat{b}_{-1}.$ (51)
According to Remark 1, for Eqs. (48), (49) and (III.2), Eq. (III.1) has the following form
$A_{00}=a_{0}+a_{1}+a_{-1}+\hat{a}_{0}+\hat{a}_{1}+\hat{a}_{-1}=1,~{}A_{10}=-(\hat{a}_{0}+\hat{a}_{1}+\hat{a}_{-1})-1,~{}A_{11}=a_{1}-a_{-1}+\hat{a}_{1}-\hat{a}_{-1},$
$A_{20}=\frac{1}{2}(\hat{a}_{0}+\hat{a}_{1}+\hat{a}_{-1}-1),~{}A_{21}=-(\hat{a}_{1}-\hat{a}_{-1}),~{}A_{22}=\frac{1}{2}(a_{1}+a_{-1}+\hat{a}_{1}+\hat{a}_{-1}),$ (52a)
$B_{00}=b_{0}+b_{1}+b_{-1}+\hat{b}_{0}+\hat{b}_{1}+\hat{b}_{-1},~{}B_{10}=-(\hat{b}_{0}+\hat{b}_{1}+\hat{b}_{-1}),B_{11}=b_{1}-b_{-1}+\hat{b}_{1}-\hat{b}_{-1},$ (52b)
where we can obtain the relation $A_{20}=-A_{10}/2-1$.
We note that Eq. (49) or (III.2) is a general form of MMFD method with a three-level structure, and is also a unified framework of FD method for convection-diffusion-like equations in $d$ dimensional space. Some existed FD schemes derived from the LBM can be taken as its special cases through selecting the parameters $a_{k},\hat{a}_{k},b_{k}$ and $\hat{b}_{k}$.
(1) SRT LB model with $\tau=1$. Set $a_{0}=a_{1}=0,a_{-1}=1;\hat{a}_{0}=\hat{a}_{1}=\hat{a}_{-1}=0$. $b_{0}=b_{1}=\hat{b}_{0}=\hat{b}_{1}=0,b_{-1}=3/2,\hat{b}_{-1}=-1/2$.
(2) TRT LB model with $s^{-}+s^{+}=2$[9, 10, 22]. Set $a_{0}=0,a_{1}=1-s^{-},a_{-1}=1;\hat{a}_{0}=-(1-s^{-}),\hat{a}_{1}=\hat{a}_{-1}=0$. $b_{0}=\hat{b}_{0}=-(1-s^{-})/2,b_{1}=\hat{b}_{1}=0,b_{-1}=3/2,\hat{b}_{-1}=-1/2$.
(3) One-stage simplified LB model without source term[25, 26]. Set $a_{0}=2(1-\tau),a_{1}=\tau-1,a_{-1}=\tau;\hat{a}_{0}=\hat{a}_{1}=\hat{a}_{-1}=0$.
(4) A new FD scheme for the wave equation without source term (Eq. (45)). Set $a_{0}=2(1-\gamma),a_{1}=a_{-1}=\gamma;\hat{a}_{0}=-1,\hat{a}_{1}=\hat{a}_{-1}=0$, where $\gamma=\alpha/(\beta c_{s}^{2})$.
In general, to derive a FD scheme one can determine the parameters with Eq. (52). We now show that Eq. (52a) is a solvable linear system. In fact, let $X=(a_{1},a_{-1},\hat{a}_{0},\hat{a}_{-1},\hat{a}_{-1})^{T}$, $b=(1-a_{0},1+A_{10},A_{11},A_{21},A_{22})^{T}$, then Eq.(52a) can be transformed into a solvable linear system $AX=b$ with the invertible coefficient matrix $A$,
$A=\left(\begin{array}[]{ccccc}1&1&1&1&1\\ 0&0&-1&-1&-1\\ 1&-1&0&1&-1\\ 0&0&0&-1&1\\ 1/2&1/2&0&1/2&1/2\\ \end{array}\right),$ (53)
and the inverse $A^{-1}$ can be obtained
$A^{-1}=\frac{1}{2}\left(\begin{array}[]{ccccc}1&1&1&1&0\\ 1&1&-1&-1&0\\ 2&0&0&0&-4\\ -1&-1&0&-1&2\\ -1&-1&0&1&2\\ \end{array}\right),$ (54)
therefore, we have
$\left(\begin{array}[]{c}a_{1}\\ a_{-1}\\ \hat{a}_{0}\\ \hat{a}_{1}\\ \hat{a}_{-1}\\ \end{array}\right)=X=A^{-1}b=\frac{1}{2}\left(\begin{array}[]{c}-a_{0}+A_{10}+A_{11}+A_{21}+2\\ -a_{0}+A_{10}-A_{11}-A_{21}+2\\ -2a_{0}-4A_{22}+2\\ a_{0}-A_{10}-A_{21}+2A_{22}-2\\ a_{0}-A_{10}+A_{21}+2A_{22}-2\\ \end{array}\right),$ (55)
with $A_{20}=-A_{10}/2-1$, while $\{b_{k},\hat{b}_{k},k=-1,0,1\}$ are of multiple selection.
III.3 The stability of mesoscopic numerical method based macroscopic finite difference scheme for linear CDE
In this part, we consider the stability of MMFD scheme (4) for the linear CDE (LCDE) without the source term. The LCDE has the following form,
$\partial_{t}\phi+\nabla\cdot\mathbf{u}\phi=\nabla\cdot\big{[}\alpha\nabla\cdot\mathbf{D}_{0}\phi\big{]},$ (56)
where $\mathbf{u}$ is a constant vector, and $\mathbf{D}_{0}$ is a constant matrix (tensor). When $\mathbf{D}_{0}=\mathbf{0}$ the LCDE above becomes a CE. The MesoFD scheme for Eq. (56) can be simplified as
$f_{j}(\mathbf{x},t)=\sum_{k,q}a_{kq}f_{j}^{eq[k,q]}.$ (57)
Let $\hat{f}(t)$ be the Fourier transformation of a function $f(\mathbf{x},t)$, then we have
$\hat{f}_{j}(t)=\sum_{k,q}a_{kq}E_{j}^{k}\hat{f}_{j}^{eq}(t+q\Delta t),E_{j}=\exp[\textbf{i}\xi_{j}],\xi_{j}=\Delta t\mathbf{\mathbf{c}}_{j}\cdot\xi.$ (58)
For Eq. (56), $f^{eq}_{j}$ has the following form [17]
$f_{j}^{eq}=\phi\bar{u}_{j},\bar{u}_{j}=\omega_{j}\left(1+\frac{\mathbf{c}_{j}\cdot\mathbf{u}}{c_{s}^{2}}+\frac{\mathbf{C}_{\alpha\alpha}\mathbf{Q}_{j\alpha\alpha}}{c_{s}^{2}(c_{\alpha}^{2}-c_{s}^{2})}+\frac{\mathbf{C}_{\alpha\bar{\alpha}}\mathbf{Q}_{j\alpha\bar{\alpha}}}{2c_{s}^{4}}\right),$ (59)
where the $\omega_{j}$ is the weight coefficient (see Appendix A for details), and $\mathbf{C}=c_{s}^{2}(\mathbf{\beta D}_{0}-\mathbf{I})+\bar{\lambda}\mathbf{uu},~{}\mathbf{Q}_{j}=\mathbf{c}_{j}\mathbf{c}_{j}-c_{s}^{2}\mathbf{I},~{}c_{s}^{2}=d_{0\alpha}c_{\alpha}^{2},~{}c_{\alpha}=\Delta x_{\alpha}/\Delta t$, and $\bar{\lambda}=0$ or 1, then summing Eq. (58), we obtain
$\hat{\phi}(t)=\sum_{k,q}a_{kq}\sum_{j}E_{j}^{k}\bar{u}_{j}\hat{\phi}(t+q\Delta t).$ (60)
Let
$\bar{u}_{j}=\bar{u}_{j}^{+}+\bar{u}_{j}^{-},~{}\bar{u}_{j}^{+}=\omega_{j}\left(1+\frac{\mathbf{C}_{\alpha\alpha}\mathbf{Q}_{j\alpha\alpha}}{c_{s}^{2}(c_{\alpha}^{2}-c_{s}^{2})}+\frac{\mathbf{C}_{\alpha\bar{\alpha}}\mathbf{Q}_{j\alpha\bar{\alpha}}}{2c_{s}^{4}}\right),~{}\bar{u}_{j}^{-}=\omega_{j}\frac{\mathbf{c}_{j}\cdot\mathbf{u}}{c_{s}^{2}},~{}A_{q}=\sum_{k}a_{kq}\sum_{j}E_{j}^{k}\bar{u}_{j},$ (61)
we can obtain
$\hat{\phi}(t)=\sum_{q}A_{q}\hat{\phi}(t+q\Delta t),A_{q}=\sum_{k}a_{kq}\sum_{j\geq 0}\big{[}\bar{u}_{j}^{+}\cos(k\xi_{j})+\textbf{i}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]},~{}q\leq 0,$ (62)
or equivalently
$\hat{\phi}(t)=\sum_{q<0}\frac{A_{q}}{1-A_{0}}\hat{\phi}(t+q\Delta t),~{}A_{q}=\sum_{k}a_{kq}\sum_{j\geq 0}\big{[}\bar{u}_{j}^{+}\cos(k\xi_{j})+\textbf{i}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]},~{}q\leq 0.$ (63)
For a two-level scheme with $q\in\{0,-1\}$, Eq. (63) is simplified as
$\hat{\phi}(t)=\frac{A_{-1}}{1-A_{0}}\hat{\phi}(t-\Delta t),A_{q}=\sum_{k}a_{kq}\sum_{j\geq 0}\big{[}\bar{u}_{j}^{+}\cos(k\xi_{j})+\textbf{i}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]},~{}q=0,-1,$ (64)
with the stability condition
$|\lambda|=|\frac{A_{-1}}{1-A_{0}}|\leq 1.$ (65)
If $A_{0}=0$, then a two-level explicit scheme is obtained with
$\lambda=A_{-1}=\sum_{k=-m}^{m}a_{k}\sum_{j\geq 0}\big{[}\bar{u}_{j}^{+}\cos(k\xi_{j})+\textbf{i}\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]}$
$=1+\sum_{j\geq 1}\sum_{k=1}^{m}\big{[}(a_{k}+a_{-k})\bar{u}_{j}^{+}(\cos(k\xi_{j})-1)+\textbf{i}(a_{k}-a_{-k})\bar{u}_{j}^{-}\sin(k\xi_{j})\big{]},$ (66)
where $a_{k}=a_{k,-1},a_{-k}=a_{-k,-1},k=1,...,m$, and $m=m_{-1}$.
Note that Eq. (III.3) has the same form as Eq. (77) in Ref. [40]. Based on Eq. (III.3) the stability theorem in Ref. [40] can be extended as follows.
Theorem 1 The two-level explicit scheme is stable ($|\lambda|\leq 1$) if and only if the following conditions holds
$(a_{k}+a_{-k})\bar{u}_{j}^{+}>0,\forall k\geq 1,j\geq 1,$ (67a)
$\sum_{j\geq 1}\sum_{k=1}^{m}(a_{k}+a_{-k})\bar{u}_{j}^{+}\leq 1,$ (67b)
$\sum_{j\geq 1}\sum_{k=1}^{m}\frac{[(a_{k}-a_{-k})\bar{u}_{j}^{-}]^{2}}{(a_{k}+a_{-k})\bar{u}_{j}^{+}}\leq 1.$ (67c)
Since
$A_{00}=a_{0}+\sum_{k=1}^{m}(a_{k}+a_{-k})=1,~{}\sum_{j}\bar{u}_{j}^{+}=1,~{}\sum_{j\geq 0}\big{[}a_{0}+\sum_{k=1}^{m}(a_{k}+a_{-k})\big{]}\bar{u}_{j}^{+}=1,$ (68)
we have
$\sum_{j\geq 1}\sum_{k=1}^{m}(a_{k}+a_{-k})\bar{u}_{j}^{+}=1-(a_{0}+\bar{u}_{0}^{+}(1-a_{0}))\leq 1\Leftrightarrow a_{0}+\bar{u}_{0}^{+}(1-a_{0})\geq 0$
$\Leftrightarrow(1-a_{0})(1-\bar{u}_{0}^{+})\leq 1\Leftrightarrow(1-\bar{u}_{0}^{+})\sum_{k=1}^{m}(a_{k}+a_{-k})\leq 1.$ (69)
Eq. (67b) is equivalent to
$a_{0}+\bar{u}_{0}^{+}(1-a_{0})\geq 0.$ (70)
For the two level $\theta$-scheme (Eq. (II.2)) without the source term, one can obtain
$\lambda=\frac{1+\theta(A+iB)}{1-(1-\theta)(A+iB)},$ (71)
where
$A=\sum_{j}\sum_{k=1}^{m_{0}}(a_{k}+a_{-k})\bar{u}_{j}^{+}(\cos(k\xi_{j})-1),B=\sum_{j}\sum_{k=1}^{m_{0}}(a_{k}-a_{-k})\bar{u}_{j}^{-}\sin(k\xi_{j}),$ (72)
and $\sum_{k}a_{k}=1$ is used. Then we obtain the following equation from Eqs. (71) and (72)
$|\lambda|^{2}-1=\frac{2A+(2\theta-1)(A^{2}+B^{2})}{|1-(1-\theta)(A+iB)|^{2}}.$ (73)
Therefore, if Eq. (67a) holds, we can prove that $|\lambda|\leq 1$ for $20\leq\theta\leq 1/$ 2 due to $A\leq 0$, while the conditional stability is obtained for $\theta>1/2$. Based above analysis, we have the stability condition of the two level $\theta$-scheme (Eq. (II.2)) without the source term.
Theorem 2 The two level $\theta$-scheme (Eq. (II.2)) without the source term is (1) unconditional stable ($|\lambda|\leq 1$) for $20\leq\theta\leq 1/$ 2, (2) conditional stable for $\theta>1/2$ if Eq. (67a) holds.
Taking $\mathbf{D}_{0}=\mathbf{I},\beta=1,\bar{\lambda}=0$ in Eq. (59), the linear EDF $f_{j}^{eq}$ with $\bar{u}_{j}^{+}=\omega_{j}>0$ and $\bar{u}_{j}^{-}=\omega_{j}\frac{\mathbf{c}_{j}\cdot\mathbf{u}}{c_{s}^{2}}$ is obtained, and
$\sum_{j\geq 1}\frac{(\bar{u}_{j}^{-})^{2}}{\bar{u}_{j}^{+}}=\sum_{j\geq 1}\omega_{j}\frac{(\mathbf{c}_{j}\cdot\mathbf{\mathbf{u}})^{2}}{c_{s}^{4}}=\frac{1}{c_{s}^{4}}\sum_{j\geq 1}\omega_{j}\mathbf{c}_{j}\mathbf{c}_{j}:\mathbf{uu}=\frac{|\mathbf{u}|^{2}}{c_{s}^{2}},$ (74)
where $\sum_{j}\omega_{j}\mathbf{c}_{j}\mathbf{c}_{j}=c_{s}^{2}\mathbf{I}$ is used. From Eq. (74) it follows that Eq. (67) and Eq. (70) become
$(a_{k}+a_{-k})>0,\forall k\geq 1,$ (75a)
$(1-\omega_{0})(1-a_{0})=(1-\omega_{0})\sum_{k=1}^{m}(a_{k}+a_{-k})\leq 1,$ (75b)
$\frac{|\mathbf{u}|^{2}}{c_{s}^{2}}\sum_{k=1}^{m}\frac{(a_{k}-a_{-k})^{2}\emph{}}{(a_{k}+a_{-k})}\leq 1,$ (75c)
and
$a_{0}+\omega_{0}(1-a_{0})\geq 0.$ (76)
Further the equivalent form of Eq. (57) can be derived from Eq. (16) as
$\phi(\mathbf{x},t)=\sum_{k,q}a_{kq}\phi(\mathbf{x},t+q\Delta t)+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}(a_{kq}+a_{-kq})\bar{u}_{j}^{+}\Delta_{j,k}^{(2)}\phi(\mathbf{x},t+q\Delta t)$
$+\frac{1}{2}\sum_{q}\sum_{j}\sum_{k=1}^{m_{q}}(a_{kq}-a_{-kq})\bar{u}_{j}^{-}\Delta_{j,k}^{(1)}\phi(\mathbf{x},t+q\Delta t).$ (77)
For the two-level explicit scheme, Eq. (III.3) is simplified as
$\phi(\mathbf{x},t)=\phi(\mathbf{x},t-\Delta t)+\frac{1}{2}\sum_{j}\sum_{k=1}^{m}\big{[}(a_{k}+a_{-k})\bar{u}_{j}^{+}\Delta_{j,k}^{(2)}+(a_{k}-a_{-k})\bar{u}_{j}^{-}\Delta_{j,k}^{(1)}\big{]}\phi(\mathbf{x},t-\Delta t),$ (78)
with
$A_{10}=-\sum_{k=-m}^{m}a_{k}=-A_{00}=-1,A_{11}=\sum_{k=1}^{m}(a_{k}-a_{-k})=A_{10}=-1,A_{22}=\frac{1}{2}\sum_{k=1}^{m}(a_{k}+a_{-k}),$ (79a)
$\alpha=-\frac{A_{22}}{A_{10}}c_{s}^{2}\Delta t=\frac{c_{s}^{2}\Delta t}{2}\sum_{k=1}^{m}(a_{k}+a_{-k})\leq\frac{c_{s}^{2}\Delta t}{2(1-\omega_{0})},$ (79b)
where Eq. (III.1), Eq. (40) and Eq. (75b) are used. Note that Eq. (79b) is equivalent to Eq. (75b) or (76). When $m=1$, Eq. (79a) is simplified as
$A_{00}=a_{0}+a_{1}+a_{-1}=-A_{10}=1,A_{10}=A_{11}=a_{1}-a_{-1}=-1,A_{22}=\frac{1}{2}(a_{1}+a_{-1})=\frac{\alpha}{\Delta tc_{s}^{2}},$ (80)
and the stability condition Eq. (75) becomes
$a_{1}+a_{-1}=1-a_{0}=\frac{2\alpha}{\Delta tc_{s}^{2}}>0,$ (81a)
$(1-\omega_{0})(1-a_{0})=(1-\omega_{0})(a_{1}+a_{-1})=(1-\omega_{0})\frac{2\alpha}{\Delta tc_{s}^{2}}\leq 1,$ (81b)
$\frac{|\mathbf{u}|^{2}}{c_{s}^{2}}\leq a_{1}+a_{-1}=\frac{2\alpha}{\Delta tc_{s}^{2}},$ (81c)
then one can obtain from Eq. (80) that
$a_{0}=1-\frac{2\alpha}{\Delta tc_{s}^{2}},~{}a_{1}=\frac{1}{2}(\frac{2\alpha}{\Delta tc_{s}^{2}}-1),~{}a_{-1}=\frac{1}{2}(\frac{2\alpha}{\Delta tc_{s}^{2}}+1),$ (82)
and the stability condition Eq. (81) is equivalent to
$|\mathbf{u}|^{2}\leq\frac{2\alpha}{\Delta t}\leq\frac{c_{s}^{2}}{1-\omega_{0}}.$ (83)
[…]
$\phi(\mathbf{x},t)=\phi(\mathbf{x},t-\Delta t)+\sum_{j=1}^{\bar{m}}\sum_{k=1}^{m}\big{[}(a_{k}+a_{-k})\bar{u}_{j}^{+}\Delta_{j,k}^{(2)}+(a_{k}-a_{-k})\bar{u}_{j}^{-}\Delta_{j,k}^{(1)}\big{]}\phi(\mathbf{x},t-\Delta t).$ (84)
When $m=1$, Eq. (84) becomes
$\phi(\mathbf{x},t)$ $=$ $\phi(\mathbf{x},t-\Delta t)+\sum_{j=1}^{\bar{m}}\big{[}(a_{1}+a_{-1})\bar{u}_{j}^{+}\Delta_{j}^{(2)}+(a_{1}-a_{-1})\bar{u}_{j}^{-}\Delta_{j}^{(1)}\big{]}\phi(\mathbf{x},t-\Delta t)$ (85)
$=$ $\phi(\mathbf{x},t-\Delta t)+\sum_{j=1}^{\bar{m}}\left[\frac{2\alpha\bar{u}_{j}^{+}}{\Delta tc_{s}^{2}}\Delta_{j}^{(2)}-\bar{u}_{j}^{-}\Delta_{j}^{(1)}\right]\phi(\mathbf{x},t-\Delta t).$
[…]
$\phi(\mathbf{x},t)=\phi(\mathbf{x},t-\Delta t)+\sum_{j=1}^{d}\left[\frac{\alpha\Delta t}{\Delta x_{j}^{2}}(\phi_{j}-2\phi+\phi_{\bar{j}})-\frac{u_{j}\Delta t}{2\Delta x_{j}}(\phi_{j}-\phi_{\bar{j}})\right],$ (86a)
$\sum_{j=1}^{d}\frac{2\alpha\Delta t}{\Delta x_{j}^{2}}\leq\sum_{j=1}^{d}\frac{d_{0j}}{1-\omega_{0}}=\sum_{j=1}^{d}\frac{d_{0j}}{d_{01}+...+d_{0d}}=1,$ (86b)
[…]
$\sum_{j=1}^{d}\frac{2\alpha\Delta t}{\Delta x_{j}^{2}}\leq\sum_{j=1}^{d}\frac{d_{0}}{1-\omega_{0}}=d.$ (87)
One can find that the condition in Eq. (86b) is much stronger than that in Eq. (87) for large $d$.
III.4 The derivation of the moment equations for NSEs using DTE method
Now using the same way as above, we can derive the moment equation for NSEs from the MesoFD method (1). The basic moments of $f_{j}$, $f_{j}^{eq}$, and $F_{j}$ are given as follows,
$M_{0}=\sum_{j}f_{j}^{eq}=\sum_{j}f_{j},\ \ \mathbf{M}_{1}=\sum_{j}\mathbf{c}_{j}f_{j}^{eq}=\sum_{j}\mathbf{c}_{j}f_{j},$ (88a)
$\mathbf{M}_{2}=\sum_{j}\mathbf{c}_{j}\mathbf{c}_{j}f_{j}^{eq},\mathbf{M}_{3}=\sum_{j}\mathbf{c}_{j}\mathbf{c}_{j}\mathbf{c}_{j}f_{j}^{eq},\mathbf{M}_{k}=\sum_{j}\mathbf{c}_{j}\cdot\cdot\cdot\mathbf{c}_{j}f_{j}^{eq},~{}k>3$ (88b)
$M_{0F}=\sum_{j}F_{j},\ \ \mathbf{M}_{1F}=\sum_{j}\mathbf{c}_{j}F_{j},\ \ \mathbf{M}_{2F}=\sum_{j}\mathbf{c}_{j}\mathbf{c}_{j}F_{j},\mathbf{M}_{kF}=\sum_{j}\mathbf{c}_{j}\cdot\cdot\cdot\mathbf{c}_{j}F_{j},~{}k>2$ (88c)
where $\mathbf{M}_{k}$ and $\mathbf{M}_{kF}$ ($k\geq 0$) are the $k$-th moments of $f_{j},f_{j}^{eq}$ and $F_{j}$, respectively. Additionally, Eq. (88a) gives the following moments of nonequilibrium,
$M_{0}^{ne}=\sum_{j}f_{j}^{ne}=0,\ \ \mathbf{M}_{1}^{ne}=\sum_{j}\mathbf{c}_{j}f_{j}^{ne}=0.$ (89)
Similarly, using DTE method we can derive the moment equations based on Eq.(1) or Eq. (5). Before doing that, we first use the double distribution based scheme Eq. (5) in the form of moments:
$M_{0}(\mathbf{x},t)=\sum_{j}\sum_{k,q}\big{[}a_{kq}f_{j}^{eq[k,q]}+\Delta tb_{kq}F_{j}^{[k,q]}\big{]},$ (90a)
$\mathbf{M}_{1}(\mathbf{x},t)=\sum_{j}\sum_{k,q}\mathbf{c}_{j}\big{[}\bar{a}_{kq}f_{j}^{eq[k,q]}+\Delta t\bar{b}_{kq}F_{j}^{[k,q]}\big{]},$ (90b)
in which two different set of parameters, $\{a_{kq},b_{kq}\}$ and $\{\bar{a}_{kq},\bar{b}_{kq}\}$ are used.
Using Taylor expansion to Eq.(90) we have
$M_{0}=A_{00}M_{0}+\sum_{l=1}^{N}\Delta t^{l}\sum_{m=0}^{l}\big{[}A_{lm}(\partial_{t}^{l-m}\nabla^{m}\otimes\mathbf{M}_{m})\big{]}+\sum_{l=0}^{N-1}\Delta t^{l+1}\sum_{m=0}^{l}\big{[}B_{lm}(\partial_{t}^{l-m}\nabla^{m}\otimes\mathbf{M}_{mF})\big{]}+O(\Delta t^{N+1}),$ (91a)
$\mathbf{M}_{1}=\bar{A}_{00}\mathbf{M}_{1}+\sum_{l=1}^{N}\Delta t^{l}\sum_{m=0}^{l}\big{[}\bar{A}_{lm}(\partial_{t}^{l-m}\nabla^{m}\otimes\mathbf{M}_{m+1})\big{]}+\sum_{l=0}^{N-1}\Delta t^{l+1}\sum_{m=0}^{l}\big{[}\bar{B}_{lm}(\partial_{t}^{l-m}\nabla^{m}\otimes\mathbf{M}_{{m+1}F})\big{]}+O(\Delta t^{N+1}),N\geq 1,$ (91b)
where
$\bar{A}_{lm}=\frac{1}{l!}C_{l}^{m}\sum_{k,q}\bar{a}_{kq}q^{l-m}k^{m},\bar{B}_{lm}=\frac{1}{l!}C_{l}^{m}\sum_{k,q}\bar{b}_{kq}q^{l-m}k^{m},$ (92)
and $A_{lm}$ is defined as in Eq. (22).
Taking $\bar{A}_{00}=A_{00}=1$ in Eq. (91), we obtain a general moment equation for NSEs
$\sum_{l=1}^{N}\Delta t^{l-1}\sum_{m=0}^{l}\big{[}A_{lm}(\partial_{t}^{l-m}\nabla^{m}\otimes\mathbf{M}_{m})\big{]}+\sum_{l=0}^{N-1}\Delta t^{l}\sum_{m=0}^{l}\big{[}B_{lm}(\partial_{t}^{l-m}\nabla^{m}\otimes\mathbf{M}_{mF})\big{]}=O(\Delta t^{N}),$ (93a)
$\sum_{l=1}^{N}\Delta t^{l-1}\sum_{m=0}^{l}\big{[}\bar{A}_{lm}(\partial_{t}^{l-m}\nabla^{m}\otimes\mathbf{M}_{m+1})\big{]}+\sum_{l=0}^{N-1}\Delta t^{l}\sum_{m=0}^{l}\big{[}\bar{B}_{lm}(\partial_{t}^{l-m}\nabla^{m}\otimes\mathbf{M}_{{m+1}F})\big{]}+O(\Delta t^{N}),N\geq 1.$ (93b)
Note that Eq. (93a) is same as Eq. (30). In what follows, we only discuss Eq.(93b) which gives for $N=1,2$ that
$\bar{A}_{10}\partial_{t}\mathbf{M}_{1}+\bar{A}_{11}\nabla\cdot\mathbf{M}_{2}=-\bar{B}_{00}\mathbf{M}_{1F}+O(\Delta t),$ (94a)
$\bar{A}_{10}\partial_{t}\mathbf{M}_{1}+\bar{A}_{11}\nabla\cdot\mathbf{M}_{2}+\Delta t\big{[}\bar{A}_{20}\partial_{t}^{2}\mathbf{M}_{1}+\bar{A}_{21}\partial_{t}\nabla\cdot\mathbf{M}_{2}+\bar{A}_{22}\nabla\nabla:\mathbf{M}_{3}\big{]}$
$=-\bar{B}_{00}\mathbf{M}_{1F}-\Delta t(\bar{B}_{10}\partial_{t}\mathbf{M}_{1F}+\bar{B}_{11}\nabla\cdot\mathbf{M}_{2F})+O(\Delta t^{2}).$ (94b)
Taking
$A_{10}=A_{11}=-B_{00},-\frac{B_{00}}{A_{10}}M_{0F}=S,$ (95a)
$\bar{A}_{10}=\bar{A}_{11}=-\bar{B}_{00},-\frac{\bar{B}_{00}}{\bar{A}_{10}}\mathbf{M}_{1F}=\mathbf{F},$ (95b)
Eq. (94) becomes
$\partial_{t}\mathbf{M}_{1}+\nabla\cdot\mathbf{M}_{2}=\mathbf{F}+O(\Delta t),$ (96a)
$\partial_{t}\mathbf{M}_{1}+\nabla\cdot\mathbf{M}_{2}+\frac{\Delta t}{\bar{A}_{10}}\big{[}\bar{A}_{20}\partial_{t}\mathbf{F}+\bar{B}_{10}\partial_{t}\mathbf{F}\big{]}$
$=\mathbf{F}-\frac{\Delta t}{\bar{A}_{10}}\big{[}(\bar{A}_{21}-\bar{A}_{20})\partial_{t}\nabla\cdot\mathbf{M}_{2}+\bar{B}_{11}\nabla\cdot\mathbf{M}_{2F}+\bar{A}_{22}\nabla\nabla:\mathbf{M}_{3}\big{]}+O(\Delta t^{2}),$ (96b)
where
$\partial_{t}^{2}\mathbf{M}_{1}=-\partial_{t}\nabla\cdot\mathbf{M}_{2}+\partial_{t}\mathbf{F}+O(\Delta t)$ (97)
is used. If we set
$\bar{A}_{20}\partial_{t}\mathbf{F}+\bar{B}_{10}\partial_{t}\mathbf{F}=(\bar{A}_{20}+\bar{B}_{10})\partial_{t}\mathbf{F}=0,$ (98)
the moment equation in the leading order of $O(\Delta t^{2})$ can be derived from Eq. (96b),
$\partial_{t}\mathbf{M}_{1}+\nabla\cdot\mathbf{M}_{2}=\mathbf{F}-\frac{\Delta t}{\bar{A}_{10}}\big{[}(\bar{A}_{21}-\bar{A}_{20})\partial_{t}\nabla\cdot\mathbf{M}_{2}+\bar{B}_{11}\nabla\cdot\mathbf{M}_{2F}+\bar{A}_{22}\nabla\nabla:\mathbf{M}_{3}\big{]}+O(\Delta t^{2}).$ (99)
Using Eq. (96a), and setting $-B_{11}\mathbf{F}=(A_{21}-A_{20})\mathbf{F}$, Eq. (39) becomes
$\partial_{t}M_{0}+\nabla\cdot\mathbf{M}_{1}=S-\frac{\Delta t}{A_{10}}\big{[}(A_{22}-(A_{21}-A_{20}))\nabla\nabla:\mathbf{M}_{2}\big{]}+O(\Delta t^{2}).$ (100)
Further according to a given NSEs, and choosing the moments and parameters in Eqs. (39) and (99) properly, we can correctly recover the NSEs from Eqs. (39) (or (100)) and (99). For example, taking
$M_{0}=\rho,\mathbf{M}_{1}=\rho\mathbf{u},\mathbf{M}_{2}=\rho\mathbf{uu}+p\mathbf{I},\mathbf{M}_{3}=c_{s}^{2}\mathbf{\Delta}\cdot\rho\mathbf{u}+\mathbf{M}_{30},$ (101a)
$\mathbf{M}_{2F}=-\frac{1}{\bar{B}_{11}}\big{[}(\bar{A}_{21}-\bar{A}_{20})\partial_{t}\mathbf{M}_{2}+\bar{A}_{22}(c_{s}^{2}(\nabla\cdot\rho\mathbf{uI}+\mathbf{u}\nabla\rho+(\mathbf{u}\nabla\rho)^{T})+\nabla\cdot\mathbf{M}_{30})\big{]},$ (101b)
$\nu=-\frac{\bar{A}_{22}}{\bar{A}_{10}}\Delta tc_{s}^{2}$ (101c)
in Eqs. (100) and (99) leads to the general NSEs
$\partial_{t}\rho+\nabla\cdot\rho\mathbf{u}=S-\frac{\Delta t}{A_{10}}\nabla\cdot\big{[}(A_{22}-(A_{21}-A_{20}))\nabla\cdot(\rho\mathbf{uu}+p\mathbf{I})\big{]}+O(\Delta t^{2}),$ (102a)
$\partial_{t}\rho\mathbf{u}+\nabla\cdot(\rho\mathbf{uu}+p\mathbf{I})=\mathbf{F}+\nu\nabla\cdot\rho\big{[}\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}\big{]}+O(\Delta t^{2}),$ (102b)
where the auxiliary source term $\mathbf{M}_{30}$ contains the additional term caused by the anisotropy of lattice tensor in a given lattice model as mentioned in Ref. [17]. If an isotropic lattice is used, $\mathbf{M}_{30}=\mathbf{0}$ can be taken.
For the incompressible flow, $\nabla\cdot(\rho\mathbf{uu}+p\mathbf{I})=O(Ma^{2})$, thus Eq. (102(a)) becomes
$\partial_{t}\rho+\nabla\cdot\rho\mathbf{u}=S+O(\Delta t^{2}+\Delta tMa^{2}),$ (103)
therefore, the incompressible NSEs are recovered from Eqs. (103) and (102(b)) in the weak compressible limit, and $\mathbf{M}_{2F}$ can be simplified [17].
III.5 The equilibrium, and source distribution functions of MesoDF method for NCDE and NSEs
[…]
$g_{j}^{eq}(A,\mathbf{B},\mathbf{C})=\omega_{j}\left(A+\tilde{\textbf{c}}_{j}\cdot{\mathbf{B}}+\tilde{\textbf{Q}}_{j}:\mathbf{C}\right),$ (104)
where $\omega_{j}$ is the weight coefficients (see Appendix A), and
$\mathbf{\tilde{c}}_{j\alpha}=\mathbf{c}_{j\alpha}/c_{s\alpha}^{2},\mathbf{\tilde{Q}}_{j\alpha\alpha}=\mathbf{Q}_{j\alpha\alpha}/(c_{s\alpha}^{2}(c_{\alpha}^{2}-c_{s\alpha}^{2})),\mathbf{\tilde{Q}}_{j\alpha\beta}=\mathbf{Q}_{j\alpha\beta}/(2c_{s\alpha}^{2}c_{s\beta}^{2})\ \ (\alpha\neq\beta),\mathbf{Q}_{j}=\mathbf{c}_{j}\mathbf{c}_{j}-\mathbf{\Delta}^{(2)}.$ (105)
Here $\mathbf{\Delta}^{(2)}=\sum_{j}\omega_{j}\mathbf{c}_{j}\mathbf{c}_{j}=\mathbf{diag}\{c_{s\alpha}\}$. We would also like to point out that strictly speaking, $c_{s\alpha}$ is direction-dependent, while it is usually taken to be a direction-independent form with $c_{s\alpha}=c_{s}$, $c_{s}$ is an adjustable parameter [12, 7, 17].
[…]
$\sum_{j}g_{j}^{eq}=A,\ \ \sum_{j}\mathbf{c}_{j}g_{j}^{eq}=\mathbf{B},\ \ \sum_{j}\mathbf{c}_{j}\mathbf{c}_{j}g_{j}^{eq}=A\mathbf{\Delta}^{(2)}+\mathbf{C},$ (106a)
$\sum_{j}\mathbf{c}_{j}\mathbf{c}_{j}\mathbf{c}_{j}g_{j}^{eq}=\mathbf{\Delta}^{(4)}\cdot\tilde{\mathbf{B}},$ (106b)
where $\tilde{\mathbf{B}}_{\alpha}=\mathbf{B}_{\alpha}/c_{s\alpha}^{2}$, and $\mathbf{\Delta}^{(4)}=\sum_{j}\omega_{j}\mathbf{c}_{j}\mathbf{c}_{j}\mathbf{c}_{j}\mathbf{c}_{j}=<\mathbf{\Delta}^{(2)}\mathbf{\Delta}^{(2)}>+\delta^{(4)}$ which is defined as
$<\mathbf{\Delta}^{(2)}\mathbf{\Delta}^{(2)}>_{\alpha\beta\gamma\theta}=\Delta_{\alpha\beta}^{(2)}\Delta_{\gamma\theta}^{(2)}+\Delta_{\alpha\gamma}^{(2)}\Delta_{\beta\theta}^{(2)}+\Delta_{\beta\gamma}^{(2)}\Delta_{\alpha\theta}^{(2)},\$ (107a)
$\mathbf{\delta}^{(4)}_{\alpha\beta\gamma\theta}=c_{s\alpha}^{2}(c_{\alpha}^{2}-3c_{s\alpha}^{2}),\alpha=\beta=\gamma=\theta;\\ \mathbf{\delta}^{(4)}_{\alpha\beta\gamma\theta}=0,else.$ (107b)
With the help of Eq. (107), we can rewrite Eq. (106b) as
$\sum_{j}\mathbf{c}_{j\alpha}\mathbf{c}_{j\beta}\mathbf{c}_{j\gamma}g_{j}^{eq}=\Delta_{\alpha\beta}^{(2)}B_{\gamma}+\Delta_{\alpha\gamma}^{(2)}B_{\beta}+\Delta_{\beta\gamma}^{(2)}B_{\alpha}+\delta^{(4)}_{\alpha\beta\gamma\theta}\tilde{B}_{\theta}.$ (108)
For the NCDE, to satisfy the moment conditions in Eqs. (34) and (40), the explicit expressions of $f_{j}^{eq}$ and $F_{j}$ with $c_{s\alpha}=c_{s}$ for all $\alpha$ can be obtained by
$f_{j}^{eq}$ $=$ $g_{j}^{eq}(\phi,\mathbf{B},\beta c_{s}^{2}\mathbf{D}+\mathbf{C}-c_{s}^{2}\phi\mathbf{I})$ (109)
$=$ $\omega_{j}\left[\phi+\frac{\textbf{c}_{j\alpha}{\mathbf{B}_{\alpha}}}{c_{s}^{2}}+\frac{(\beta c_{s}^{2}\mathbf{D}+\mathbf{C}-c_{s}^{2}\phi\mathbf{I})_{\alpha\alpha}(c_{j\alpha}^{2}-c_{s}^{2})}{c_{s}^{2}(c_{\alpha}^{2}-c_{s}^{2})}+\frac{(\beta c_{s}^{2}\mathbf{D}+\mathbf{C}-c_{s}^{2}\phi\mathbf{I})_{\alpha\bar{\alpha}}(c_{j\alpha}c_{j\bar{\alpha}})}{2c_{s}^{4}}\right],$
$F_{j}=g_{j}^{eq}(S,\mathbf{M}_{1F},\mathbf{0})=\omega_{j}(S+\frac{\mathbf{c}_{j\alpha}\mathbf{M}_{1F,\alpha}}{c_{s}^{2}}),$ (110)
where $\bar{\alpha}$ denotes the index $\gamma$ with $\gamma\neq\alpha$, $\mathbf{M}_{1F}$ is given by Eq. (40).
While for the NSEs, to satisfy the moment conditions in Eqs. (95) and (101), $f_{j}^{eq}$ and $F_{j}$ with $c_{s\alpha}=c_{s}$ for all $\alpha$ can be given by
$f_{j}^{eq}$ $=$ $g_{j}^{eq}(\rho,\rho\mathbf{u},\rho\mathbf{uu})=\rho g_{j}^{eq}(1,\mathbf{u},\mathbf{uu})$ (111)
$=$ $\omega_{j}\rho\left[1+\frac{\textbf{c}_{j\alpha}{\mathbf{u}_{\alpha}}}{c_{s}^{2}}+\frac{u_{\alpha}^{2}(c_{j\alpha}^{2}-c_{s}^{2})}{c_{s}^{2}(c_{\alpha}^{2}-c_{s}^{2})}+\frac{u_{\alpha}u_{\bar{\alpha}}(c_{j\alpha}c_{j\bar{\alpha}})}{2c_{s}^{4}}\right],$
$F_{j}=g_{j}^{eq}(S,\mathbf{F},\mathbf{M}_{2F}-c_{s}^{2}S\mathbf{I})=\omega_{j}\left[S+\frac{\textbf{c}_{j\alpha}{\mathbf{\mathbf{F}}_{\alpha}}}{c_{s}^{2}}+\frac{(\mathbf{M}_{2F,\alpha\alpha}-c_{s}^{2}S)(c_{j\alpha}^{2}-c_{s}^{2})}{c_{s}^{2}(c_{\alpha}^{2}-c_{s}^{2})}+\frac{\mathbf{M}_{2F,\alpha\bar{\alpha}}(c_{j\alpha}c_{j\bar{\alpha}})}{2c_{s}^{4}}\right],$ (112)
where $\mathbf{M}_{2F}$ is determined by Eq. (101b).
Remark 2 : For incompressible flow, inserting $\rho=\rho_{0}+\delta\rho$ with $\delta\rho=O(Ma^{2})$ into Eq. (111), and omitting the terms of $O(Ma^{3})$, one can obtain the He-Luo-like version [42] of the present MesoFD scheme with $f_{j}^{eq}=g_{j}^{eq}(\rho,\rho_{0}\mathbf{u},\rho_{0}\mathbf{uu})$.
IV Numerical tests
In this section, several different numerical examples are used to test the present MesoFD scheme. Here we only use the MMFD scheme (49). To test the accuracy of the MesoFD scheme, the following global relative errors (GRE) are adopted:
$\operatorname{GRE}=\frac{\sum_{i}\left|\phi\left(\bm{x}_{i},t\right)-\phi^{*}\left(\bm{x}_{i},t\right)\right|}{\sum_{i}\left|\phi^{*}\left(\bm{x}_{i},t\right)\right|},$
where $\phi(\bm{x},t)$ and $\phi^{*}(\bm{x},t)$ denote numerical and analytical solutions, respectively.
Example 1: Isotropic convection-diffusion equation
The following two-dimensional isotropic CDE is first considered to test the accuracy of the MesoFD scheme,
$\partial_{t}\phi+u_{x}\partial_{x}\phi+u_{y}\partial_{y}\phi=\kappa\left(\partial_{xx}\phi+\partial_{yy}\phi\right)+R,$
where $u_{x}=u_{y}=0.1,~{}\kappa$ is the diffusion coefficient, and the source term $R$ is given by
$R=\exp\left[\left(1-2\pi^{2}\kappa\right)t\right]\left(\sin[\pi(x+y)]+\pi\left(u_{x}+u_{y}\right)\cos[\pi(x+y)]\right).$
The analytical solution of this problem can be expressed as
$\phi(x,y,t)=\exp\left[\left(1-2\pi^{2}\kappa\right)t\right]\sin(\pi(x+y)),$
and we also have the initial condition
$\phi(x,y,0)=\sin(\pi(x+y)).$
The computation domain is $(x,y)\in[0,2]\times[0,2]$ and $t\in[0,T]$.
We take
$f_{j}^{eq}=\omega_{j}\phi\left[1+\frac{\mathbf{c}_{j}\cdot\mathbf{u}}{c_{s}^{2}}+\frac{(\mathbf{u}\mathbf{u}+\beta c_{s}^{2}\mathbf{I}-c_{s}^{2}\mathbf{I}):(\mathbf{c}_{j}\mathbf{c}_{j}-c_{s}^{2}\mathbf{I})}{2c_{s}^{4}}\right],$
and
$F_{j}=\omega_{j}R\left[1+\frac{\mathbf{c}_{j}\cdot\mathbf{u}}{c_{s}^{2}}\right].$
In the simulation, we use the D2Q9 lattice model with $\omega_{0}=4/9,~{}\omega_{1-4}=1/9,~{}\omega_{5-8}=1/36$, and $c_{s}^{2}=c^{2}/3,~{}c=\Delta x/\Delta t$. Then we consider the following cases:
Case 1: SRT LB model with $\tau=1$. Let $A_{10}=A_{11}=-B_{00}=-1,~{}A_{21}=0,~{}A_{22}=\frac{1}{2},~{}a_{0}=0$, then we have
$a_{-1}=1,\quad a_{0}=0,\quad a_{1}=0,\quad\hat{a}_{-1}=0,\quad\hat{a}_{0}=0,\quad\hat{a}_{1}=0,$
$b_{0}=b_{1}=\hat{b}_{0}=\hat{b}_{1}=0,\quad b_{-1}=\frac{3}{2},\quad\hat{b}_{-1}=-\frac{1}{2},$
the MMFD scheme (49) becomes
$\phi(\mathbf{x},t+\Delta t)=\sum_{j}\left(f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)+\frac{3}{2}\Delta tF_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)-\frac{1}{2}\Delta tF_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)\right).$
Case 2: Let $A_{10}=A_{11}=-B_{00}=-1,~{}A_{21}=0,~{}A_{22}=\frac{1}{2},~{}a_{0}=\frac{1}{4}$, then we have
$a_{-1}=\frac{7}{8},\quad a_{0}=\frac{1}{4},\quad a_{1}=-\frac{1}{8},\quad\hat{a}_{-1}=\frac{1}{8},\quad\hat{a}_{0}=-\frac{1}{4},\quad\hat{a}_{1}=-\frac{1}{8},$
$b_{0}=b_{1}=\hat{b}_{0}=\hat{b}_{1}=0,\quad b_{-1}=\frac{3}{2},\quad\hat{b}_{-1}=-\frac{1}{2},$
and
$\phi(\mathbf{x},t+\Delta t)$ $=\sum_{j}\left(\frac{7}{8}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)+\frac{1}{4}f_{j}^{eq}(\mathbf{x},t)-\frac{1}{8}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)\right.$
$+\frac{1}{8}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)-\frac{1}{4}f_{j}^{eq}(\mathbf{x},t-\Delta t)-\frac{1}{8}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t-\Delta t)$
$\left.+\frac{3}{2}\Delta tF_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)-\frac{1}{2}\Delta tF_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)\right).$
Case 3: Let $A_{10}=A_{11}=-B_{00}=-1,~{}A_{21}=\frac{1}{4},~{}A_{22}=\frac{1}{2},~{}a_{0}=0$, then we have
$a_{-1}=\frac{7}{8},\quad a_{0}=0,\quad a_{1}=\frac{1}{8},\quad\hat{a}_{-1}=\frac{1}{8},\quad\hat{a}_{0}=0,\quad\hat{a}_{1}=-\frac{1}{8},$
$b_{0}=b_{1}=\hat{b}_{0}=\hat{b}_{1}=0,\quad b_{-1}=\frac{3}{2},\quad\hat{b}_{-1}=-\frac{1}{2},$
and
$\phi(\mathbf{x},t+\Delta t)$ $=\sum_{j}\left(\frac{7}{8}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)+\frac{1}{8}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)+\frac{1}{8}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)-\frac{1}{8}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t-\Delta t)\right.$
$\left.+\frac{3}{2}\Delta tF_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)-\frac{1}{2}\Delta tF_{j}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)\right).$
Let $N_{x}=\frac{1}{\Delta x}$, $N_{t}=\frac{1}{\Delta t}$, then the errors and convergency orders at $T=1.0$ are listed in Table 1. From the table, the second convergency order can be found, which is in agreement with the theoretical analysis.
Table 1: The errors and convergency orders for Case 1 - 3
Case 1 Case 2 Case 3
$N_{x}$ $N_{t}$ GRE1 order1 GRE2 order2 GRE3 order3
10 6 1.51E-01 1.05E-01 1.01E-01
20 24 4.16E-02 1.86 2.50E-02 2.07 2.45E-02 2.04
40 96 1.07E-02 1.97 6.18E-03 2.01 6.08E-03 2.01
80 384 2.68E-03 1.99 1.54E-03 2.00 1.52E-03 2.00
Example 2: Isotropic convection-diffusion equation without source term
For the homogeneous isotropic convection-diffusion equation,
$\partial_{t}\phi+u_{x}\partial_{x}\phi+u_{y}\partial_{y}\phi=D\left(\partial_{xx}\phi+\partial_{yy}\phi\right),$
where $u_{x}=u_{y}=1.0$ and $D=0.02$ is the diffusion coefficient. The computation domain is $(x,y)\in[0,2]\times[0,2]$ and $t\in[0,T]$. The exact solution is
$\phi(x,y,t)=\exp\left(-2D\pi^{2}t\right)\cos[\pi(x+y)-\pi(u_{x}+u_{y})t],$
and we also have the initial condition
$\phi(x,y,0)=\cos(\pi(x+y)).$
We take
$f_{j}^{eq}=\omega_{j}\phi[1+\frac{\mathbf{c}_{j}\cdot\mathbf{u}}{c_{s}^{2}}+\frac{(\mathbf{u}\mathbf{u}+\beta c_{s}^{2}\mathbf{I}-c_{s}^{2}\mathbf{I}):(\mathbf{c}_{j}\mathbf{c}_{j}-c_{s}^{2}\mathbf{I})}{2c_{s}^{4}}],$
and use D2Q9 lattice model with $\omega_{0}=4/9,~{}\omega_{1-4}=1/9,~{}\omega_{5-8}=1/36$. Here, we also consider the following three cases:
Case 1: Let $A_{10}=A_{11}=-1,~{}A_{21}=0,~{}A_{22}=\frac{1}{2},~{}a_{0}=0,~{}\beta=1$, then we have
$a_{-1}=1,\quad a_{0}=0,\quad a_{1}=0,\quad\hat{a}_{-1}=0,\quad\hat{a}_{0}=0,\quad\hat{a}_{1}=0,$
and
$\phi(\mathbf{x},t+\Delta t)=\sum_{j}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t).$
Case 2: Let $A_{10}=A_{11}=-1,~{}A_{21}=\frac{1}{4},~{}A_{22}=\frac{1}{2},~{}a_{0}=0,~{}\beta=1$, then we have
$a_{-1}=\frac{7}{8},\quad a_{0}=0,\quad a_{1}=\frac{1}{8},\quad\hat{a}_{-1}=\frac{1}{8},\quad\hat{a}_{0}=0,\quad\hat{a}_{1}=-\frac{1}{8},$
and
$\phi(\mathbf{x},t+\Delta t)=\sum_{j}\left(\frac{7}{8}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)+\frac{1}{8}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)+\frac{1}{8}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)-\frac{1}{8}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t-\Delta t)\right).$
Case 3: Let $A_{10}=A_{11}=-1,~{}A_{21}=0,~{}A_{22}=1,~{}a_{0}=0,~{}\beta=\frac{1}{2}$, then we have
$a_{-1}=1,\quad a_{0}=a_{1}=0,\quad\hat{a}_{-1}=\frac{1}{2},\quad\hat{a}_{0}=-1\quad\hat{a}_{1}=\frac{1}{2},$
and
$\phi(\mathbf{x},t+\Delta t)=\sum_{j}\left(f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)+\frac{1}{2}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)-f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t-\Delta t)+\frac{1}{2}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t-\Delta t)\right).$
Table 2: The errors and convergency orders for Case 1 - 3
Case 1 Case 2 Case 3
$N_{x}$ $N_{t}$ GRE1 order1 GRE2 order2 GRE3 order3
10 12 5.44E-01 5.74E-01 7.39E-00
20 48 7.20E-02 2.92 1.41E-02 2.02 2.41E-02 4.94
40 192 1.45E-02 2.31 3.31E-02 2.10 6.24E-02 1.95
80 768 3.34E-03 2.12 8.05E-03 2.04 1.58E-02 1.98
In the simulations, $N_{x}=\frac{1}{\Delta x}$, $N_{t}=\frac{1}{\Delta t}$, then the errors and convergency orders at $T=1.0$ are listed in Table 2. From the table, the second convergency order also can be found, which is in agreement with the theoretical analysis. In order to test the stable condition, that is,
$\frac{\mathbf{u}^{2}}{c_{s}^{2}}\leq\frac{1}{1-\omega_{0}}=\frac{9}{5}.$
for the two-level scheme. We use case 1 to verify the stable condition, and take $D=0.01,~{}u_{x}=u_{y}=1.0$. The errors are list in Table 3. From the tabel, we can see that if $\frac{\mathbf{u}^{2}}{c_{s}^{2}}$ is more than 9/5, then the scheme is unstable which is agreement with the Theorem 1.
Table 3: The errors with different $c_{s}^{2}$
$\Delta x$ $\Delta t$ $\frac{\mathbf{u}^{2}}{c_{s}^{2}}$ GRE
$\frac{1}{10}$ $\frac{1}{6}$ $\frac{50}{3}$ 7.78E+02
$\frac{1}{20}$ $\frac{1}{24}$ $\frac{25}{6}$ 2.02E-00
$\frac{1}{40}$ $\frac{1}{96}$ $\frac{25}{24}$ 3.26E-02
$\frac{1}{80}$ $\frac{1}{384}$ $\frac{25}{96}$ 6.90E-03
Example 3: Wave equation
For the two-dimensional wave equation
$\frac{\partial^{2}\phi}{\partial t^{2}}=\frac{1}{2}(\frac{\partial^{2}\phi}{\partial x^{2}}+\frac{\partial^{2}\phi}{\partial y^{2}}),\quad(x,y)\in[0,1]\times[0,1],t\in[0,1],$
the exact solution is
$\phi(x,y,t)=e^{x+y+t}.$
The initial conditions are
$\phi(x,y,0)=e^{x+y},\quad\frac{\partial\phi(x,y,0)}{\partial t}=e^{x+y}.$
In the simulation, we take $N_{x}=N_{y}$ and the space step is $\Delta x=\Delta y=1/N_{x}$. We consider two lattice models D2Q5 and D2Q9. Based on the idea of the finite difference method, we have the approximation of $\phi(x,\Delta t)$
$\frac{2}{\Delta t}\left[\frac{\phi(\mathbf{x},\Delta t)-\phi(\mathbf{x},0)}{\Delta t}-\frac{\partial\phi(\mathbf{x},0)}{\partial t}\right]=\frac{1}{2}\sum_{j=1}^{2}\left[\frac{\phi(\mathbf{x}-\mathbf{c}_{j}\Delta t,0)-2\phi(\mathbf{x},0)+\phi(\mathbf{x}+\mathbf{c}_{j}\Delta t,0)}{\Delta x^{2}}\right],$
that is,
$\phi(\mathbf{x},\Delta t)=\phi(\mathbf{x},0)+\Delta t\frac{\partial\phi(\mathbf{x},0)}{\partial t}+\left(\frac{\Delta t}{2\Delta x}\right)^{2}\sum_{j=1}^{2}[\phi(\mathbf{x}-\mathbf{c}_{j}\Delta t,0)-2\phi(\mathbf{x},0)+\phi(\mathbf{x}+\mathbf{c}_{j}\Delta t,0)].$
We take the equilibrium function as follows
$f_{j}^{eq}=\omega_{j}\phi\left[1+\frac{(\beta c_{s}^{2}\mathbf{I}-c_{s}^{2}\mathbf{I}):(\mathbf{c}_{j}\mathbf{c}_{j}-c_{s}^{2}\mathbf{I})}{2c_{s}^{4}}\right].$
Here $a_{0}=2(1-\gamma),~{}a_{1}=a_{-1}=\gamma,~{}\hat{a}_{0}=-1,~{}\hat{a}_{1}=\hat{a}_{-1}=0$, $A_{22}=\gamma$.
Case 1: We use the D2Q5 model and let $\beta=1,\gamma=\frac{3}{8\beta}=\frac{3}{8}$, we have
$\phi(\mathbf{x},t+\Delta t)$ $=\sum_{j}\left(\frac{3}{8}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)+\frac{5}{4}f_{j}^{eq}(\mathbf{x},t)+\frac{3}{8}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)-f_{j}^{eq}(\mathbf{x},t-\Delta t)\right).$
Case 2: We use the D2Q9 model and the parameters are the same as in Case 1.
Case 3: We use the D2Q9 model and take $\beta=\frac{3}{4},~{}\gamma=\frac{3}{8\beta}=\frac{1}{2}$, we obtain
$\phi(\mathbf{x},t+\Delta t)$ $=\sum_{j}\left(\frac{1}{2}f_{j}^{eq}(\mathbf{x}-\mathbf{c}_{j}\Delta t,t)+f_{j}^{eq}(\mathbf{x},t)+\frac{1}{2}f_{j}^{eq}(\mathbf{x}+\mathbf{c}_{j}\Delta t,t)-f_{j}^{eq}(\mathbf{x},t-\Delta t)\right).$
It is noted that the scheme in Case 1 is the same as the explicit finite-difference scheme. The errors and convergency orders are listed in Table 4. The second-order convergency can be found and we can obtain some different schemes which cannot be derived from finite-difference method directly.
Table 4: The errors and convergency orders for Case 1-3 of the wave equation
Case 1 Case 2 Case 3
$N_{x}$ $N_{t}$ GRE1 order GRE2 order2 GRE3 order3
5 10 2.86E-04 6.04E-03 3.91E-03
10 20 6.09E-05 2.23 1.74E-03 1.79 1.13E-03 1.79
20 40 1.39E-05 2.13 4.55E-04 1.94 2.97E-04 1.93
40 80 3.30E-06 2.07 1.16E-04 1.98 7.58E-05 1.97
V Conclusions
By utilizing the equilibrium distribution function to represent the distribution function and/or the nonequilibrium distribution function, a new kind of the mesoscopic numerical method (MesoNM), namely the mesoscopic finite difference method (MesoFDM) with a multi-level and multi-point structure is presented for PDEs including the NSEs and NCDE. The macroscopic version of the MesoFD scheme, the MMFD scheme for the conserved variables, can be obtained directly from the moments of the MesoFD scheme.
A detailed DTE analysis on the present MesoFDM is conducted, and the macroscopic moment equations with truncation errors of different orders can be obtained. The NSEs and NCDE can be recovered from these moment equations to second-order accuracy in space. The MMFD scheme is a unified FD scheme. The existing FD schemes derived from the LB models and even the traditional FD schemes are the special cases of the present MMFD scheme. The MesoFDM not only establishes a connection between the MesoNM and the traditional macroscopic FDM, but also extends the LB method and traditional FDM. We can combine the ’bottom-up’ mesoscopic method with the ’top-down’ macroscopic method to design more effective numerical methods by using their own advantages.
Acknowledgements
This work was financially supported by the National Natural Science Foundation of China (Grants No. 12072127, No. 51836003 and No. 12202130).
Appendix A Lattice models of RMRT-LB method on rectangular lattice [17]
[…]
rD2Q9 lattice:
$\begin{split}&\{\mathbf{c}_{j},0\leq j\leq 8\}=\left(\begin{array}[]{ccccccccc}0&c_{1}&0&-c_{1}&0&c_{1}&-c_{1}&-c_{1}&c_{1}\\ 0&0&c_{2}&0&-c_{2}&c_{2}&c_{2}&-c_{2}&-c_{2}\\ \end{array}\right),\\ &\omega_{j}\geq 0,\omega_{1}=\omega_{3},\omega_{2}=\omega_{4},\omega_{5}=\omega_{6}=\omega_{7}=\omega_{8},\omega_{0}=1-2(\omega_{1}+\omega_{2}+2\omega_{5}),\\ &\omega_{0}=(1-d_{01})(1-d_{02}),\omega_{1}=d_{01}(1-d_{02})/2,\omega_{2}=(1-d_{01})d_{02}/2,\omega_{5}=d_{01}d_{02}/4,\end{split}$ (113)
where $d_{01},d_{02}\in(0,1)$.
rD2Q5I lattice:
$\begin{split}&\{\mathbf{c}_{j},0\leq j\leq 4\}=\left(\begin{array}[]{ccccc}0&c_{1}&0&-c_{1}&0\\ 0&0&c_{2}&0&-c_{2}\\ \end{array}\right),\\ &\omega_{j}\geq 0,\omega_{1}=\omega_{3},\omega_{2}=\omega_{4},\omega_{0}=1-2(\omega_{1}+\omega_{2}),\\ &\omega_{0}=1-d_{01}-d_{02},\omega_{1}=d_{01}/2,\omega_{2}=d_{02}/2,\end{split}$ (114)
where $d_{01}>0,d_{02}>0$, and $d_{01}+d_{02}<1$.
rD2Q5II lattice:
$\begin{split}&\{\mathbf{c}_{j},0\leq j\leq 4\}=\left(\begin{array}[]{ccccc}0&c_{1}&-c_{1}&-c_{1}&c_{1}\\ 0&c_{2}&c_{2}&-c_{2}&-c_{2}\\ \end{array}\right),\\ &\omega_{j}\geq 0,\omega_{1}=\omega_{2}=\omega_{3}=\omega_{4},\omega_{0}=1-4\omega_{1},\\ &\omega_{0}=1-d_{0},\omega_{1}=d_{0}/4,\end{split}$ (115)
where $d_{01}=d_{02}=d_{0}\in(0,1)$.
rD3Q27 lattice:
$\{\mathbf{c}_{j},0\leq j\leq 6\}=\left(\begin{array}[]{ccccccc}0&c_{1}&0&0&-c_{1}&0&0\\ 0&0&c_{2}&0&0&-c_{2}&0\\ 0&0&0&c_{3}&0&0&-c_{3}\\ \end{array}\right),$ (116a)
$\{\mathbf{c}_{j},7\leq j\leq 18\}=\left(\begin{array}[]{cccccccccccc}c_{1}&-c_{1}&-c_{1}&c_{1}&c_{1}&-c_{1}&-c_{1}&c_{1}&0&0&0&0\\ c_{2}&c_{2}&-c_{2}&-c_{2}&0&0&0&0&c_{2}&-c_{2}&-c_{2}&c_{2}\\ 0&0&0&0&c_{3}&-c_{3}&-c_{3}&c_{3}&c_{3}&c_{3}&-c_{3}&-c_{3}\\ \end{array}\right),$ (116b)
$\{\mathbf{c}_{j},19\leq j\leq 26\}=\left(\begin{array}[]{cccccccc}c_{1}&c_{1}&c_{1}&-c_{1}&-c_{1}&-c_{1}&-c_{1}&c_{1}\\ c_{2}&c_{2}&-c_{2}&c_{2}&-c_{2}&-c_{2}&c_{2}&-c_{2}\\ c_{3}&-c_{3}&c_{3}&c_{3}&-c_{3}&c_{3}&-c_{3}&-c_{3}\\ \end{array}\right),$ (116c)
$\begin{split}\omega_{j}\geq 0,\omega_{1}&=\omega_{4},\omega_{2}=\omega_{5},\omega_{3}=\omega_{6},\omega_{7}=\omega_{8}=\omega_{9}=\omega_{10},\omega_{11}=\omega_{12}=\omega_{13}=\omega_{14},\\ &\omega_{15}=\omega_{16}=\omega_{17}=\omega_{18},\omega_{j}=\omega_{19}(j>19),\omega_{0}=1-\sum_{j>0}\omega_{j},\end{split}$
where
$\omega_{0}=(1-d_{01})(1-d_{02})(1-d_{03}),$ (117a)
$\omega_{1}=d_{01}(1-d_{02})(1-d_{03})/2,\omega_{2}=d_{02}(1-d_{01})(1-d_{03})/2,\omega_{3}=d_{03}(1-d_{01})(1-d_{02})/2,$ (117b)
$\omega_{7}=d_{01}d_{02}(1-d_{03})/4,\omega_{11}=d_{01}d_{03}(1-d_{02})/4,\omega_{15}=d_{02}d_{03}(1-d_{01})/4,\omega_{19}=d_{01}d_{02}d_{03}/8,$ (117c)
where $d_{01},d_{02},d_{03}\in(0,1)$.
rD3Q19 lattice:
$\{\mathbf{c}_{j},0\leq j\leq 6\}=\left(\begin{array}[]{ccccccc}0&c_{1}&0&0&-c_{1}&0&0\\ 0&0&c_{2}&0&0&-c_{2}&0\\ 0&0&0&c_{3}&0&0&-c_{3}\\ \end{array}\right),$ (118a)
$\{\mathbf{c}_{j},7\leq j\leq 18\}=\left(\begin{array}[]{cccccccccccc}c_{1}&-c_{1}&-c_{1}&c_{1}&c_{1}&-c_{1}&-c_{1}&c_{1}&0&0&0&0\\ c_{2}&c_{2}&-c_{2}&-c_{2}&0&0&0&0&c_{2}&-c_{2}&-c_{2}&c_{2}\\ 0&0&0&0&c_{3}&-c_{3}&-c_{3}&c_{3}&c_{3}&c_{3}&-c_{3}&-c_{3}\\ \end{array}\right),$ (118b)
$\begin{split}\omega_{j}\geq 0,\omega_{1}&=\omega_{4},\omega_{2}=\omega_{5},\omega_{3}=\omega_{6},\omega_{7}=\omega_{8}=\omega_{9}=\omega_{10},\omega_{11}=\omega_{12}=\omega_{13}=\omega_{14},\\ &\omega_{15}=\omega_{16}=\omega_{17}=\omega_{18},\omega_{0}=1-\sum_{j>0}\omega_{j},\end{split}$
where
$\omega_{0}=1-d_{01}-d_{02}-d_{03}+d_{01}d_{02}+d_{01}d_{03}+d_{02}d_{03},$ (119a)
$\omega_{1}=d_{01}(1-d_{02}-d_{03})/2,\omega_{2}=d_{02}(1-d_{01}-d_{03})/2,\omega_{3}=d_{03}(1-d_{01}-d_{02})/2,$ (119b)
$\omega_{7}=d_{01}d_{02}/4,\omega_{11}=d_{01}d_{03}/4,\omega_{15}=d_{02}d_{03}/4,$ (119c)
where $d_{01},d_{02}$ and $d_{03}$ make the weights $\omega_{j}>0,\forall j$.
rD3Q9 lattice:
$\begin{split}&\{\mathbf{c}_{j},0\leq j\leq 8\}=\left(\begin{array}[]{ccccccccc}0&c_{1}&c_{1}&c_{1}&-c_{1}&-c_{1}&-c_{1}&-c_{1}&c_{1}\\ 0&c_{2}&c_{2}&-c_{2}&c_{2}&-c_{2}&-c_{2}&c_{2}&-c_{2}\\ 0&c_{3}&-c_{3}&c_{3}&c_{3}&-c_{3}&c_{3}&-c_{3}&-c_{3}\\ \end{array}\right),\\ &\omega_{j}\geq 0,\omega_{1}=\omega_{2}=...=\omega_{8},\omega_{0}=1-8\omega_{1},\\ &\omega_{0}=1-d_{0},\omega_{1}=d_{0}/8,\end{split}$ (120)
where $d_{01}=d_{02}=d_{03}=d_{0}\in(0,1)$
rD3Q7 lattice:
$\{\mathbf{c}_{j},0\leq j\leq 6\}=\left(\begin{array}[]{ccccccc}0&c_{1}&0&0&-c_{1}&0&0\\ 0&0&c_{2}&0&0&-c_{2}&0\\ 0&0&0&c_{3}&0&0&-c_{3}\\ \end{array}\right),$ (121a)
$\begin{split}\omega_{j}\geq 0,\omega_{1}&=\omega_{4},\omega_{2}=\omega_{5},\omega_{3}=\omega_{6},\omega_{0}=1-2(\omega_{1}+\omega_{2}+\omega_{3}),\end{split}$
where
$\omega_{0}=1-d_{01}-d_{02}-d_{03},\omega_{1}=d_{01}/2,\omega_{2}=d_{02}/2,\omega_{3}=d_{03}/2,$ (122)
where $d_{01},d_{02},d_{03}>0,d_{01}+d_{02}+d_{03}<1$
[…]